### v1

# Multidimensional Dickman distribution and operator selfdecomposability

Abstract
The one-dimensional Dickman distribution arises in various stochastic models across number theory, combinatorics, physics, and biology. Recently, a definition of the multidimensional Dickman distribution has appeared in the literature, together with its application to approximating the small jumps of multidimensional Lévy processes. In this paper, we extend this definition to a class of vector-valued random elements, which we characterise as fixed points of a specific affine transformation involving a random matrix obtained from the matrix exponential of a uniformly distributed random variable. We prove that these new distributions possess the key properties of infinite divisibility and operator selfdecomposability. Furthermore, we identify several cases where this new distribution arises as a limiting distribution.
Keywords: multivariate Dickman distribution, operator selfdecomposability, random affine transformation, random difference equation, random matrix
1 Introduction
The one-dimensional Dickman distribution, as defined in [PW04], appears as a limiting law in several settings. For instance, it arises as the long-time limit of a shot-noise process with fixed signal amplitude and exponential propagation [TAK54], and in a model of the atomic cascade process [CS75]. It can also be characterised as a fixed point of an affine transformation with random coefficients, or equivalently, as the stationary solution of a random difference equation [VER79]. Its relative, the max-Dickman distribution, initially arose in number theory [DIC30, DE 51a], and later appeared in a combinatorial setting [GON44]. Recently, a number of new applications of the Dickman distribution have been discovered in random graph theory [PW04] as well as in various limiting schemes [PIN18, CSZ19, BG19, BM20, GMP22]. The application of Dickman distribution in the theory of non-local operators was discussed in [GKL+24]. It was also shown to be useful in approximating the small jumps of positive Lévy processes in cases when Brownian approximation fails [COV09]. Some generalisations of the Dickman distribution have been considered in [ABT03, PIN18, BG19, GKL+25] For simulation from the Dickman distribution, we refer to [DF10, FH10, DQL19]. For the historical review on the Dickman distribution, see [PW04, MP20].
Recently, the generalisation of the one-dimensional Dickman distribution to the multidimensional case has appeared in [BM20] together with its application towards approximating small jumps of multidimensional Lévy processes [GZ24a]. In this paper, we propose to extend the definition given in [BM20] to a class of vector-valued random elements, which we define, in analogy to the one-dimensional version, as fixed points of a certain affine transformation with random coefficients. In particular, the affine transformation we consider is induced by translating a given point by a random vector and further multiplying it by a factor defined as the operator (matrix) exponential of a uniformly distributed random variable. We refer to the elements of this class as the operator Dickman distributions. The motivation for this extension initially comes from the property of operator selfdecomposability, which we shall demonstrate holds for members of our class. It is also supported by some applications analogous to the classical applications of the one-dimensional Dickman distribution.
The present paper is structured as follows. We recall the definition and key properties of the one-dimensional Dickman distribution in Section 2. In Section 3, we first consider the multidimensional Dickman distribution as defined in [BM20], and then introduce a broader class of operator Dickman distributions, highlighting some of their key properties. Some applications of the operator Dickman distributions are considered in Section 4. The simulations from the operator Dickman distribution are demonstrated in Section 5.
2 One-dimensional Dickman distribution
The one-dimensional Dickman distribution with parameter $\theta>0$ can be defined as a distribution of a random variable $X_{D}$ satisfying the equation
$X_{D}\overset{d}{=}U^{1/\theta}(1+X^{\prime}_{D}),$ (1)
where the symbol $\overset{d}{=}$” denotes the equality in distribution, $X^{\prime}_{D}\overset{d}{=}X_{D}$, $U$ has the uniform distribution on $[0,1]$ and is independent of $X^{\prime}_{D}$, see e.g. [PW04]. Equivalently, the random variable $X_{D}$ is the fixed point (in a distributional sense) of the random affine transformation $\cdot\mapsto U^{1/\theta}(1+\cdot)$, where $U$ is as above. We will further denote $GD_{\theta}$ the Dickman distribution with parameter $\theta>0$.
The distribution $GD_{\theta}$ is absolutely continuous for all $\theta>0$, and its density can be represented as
$f_{\theta}(x)=\frac{e^{-\gamma\theta}}{\Gamma(\theta)}\rho_{\theta}(x),\penalty 10000\ x\in\mathbb{R},$
where $\Gamma(\cdot)$ is the gamma function, $\gamma=-\Gamma^{\prime}(1)\approx 0.5772$ is Euler’s constant, and the function $\rho_{\theta}(x)$ satisfies the difference-differential equation
$x\rho^{\prime}_{\theta}(x)+(1-\theta)\rho_{\theta}(x)+\theta\rho_{\theta}(x-1)=0,\quad x>1,$ (2)
with initial conditions
$\rho_{\theta}(x)=0,\quad x\leq 0,$
$\rho_{\theta}(x)=x^{\theta-1},\quad 0<x\leq 1.$
In the case of $\theta=1$, the function $\rho_{1}(\cdot)$ is known as the Dickman function. It occurred in the work of Karl Dickman [DIC30] in the number-theoretical context from which it gained its name. However, it can be found in even earlier Ramanujan’s unpublished notes [GEO15, Ch. 8]. The equation (2) and its generalisations have been studied in [WHE90].
The density $f_{\theta}(x)$ of the $GD_{\theta}$ distribution has been an object of independent studies, and we recall here a few results related to it. The asymptotic behaviour of the density $f_{\theta}(x)$ for $\theta=1$ has been established in [DE 51b] and for general $\theta$ in [VER72, Lem. 4.7.9]. It was also proven in [VER72, Thm. 4.7.7] that the density $f_{\theta}(x)$ can be expressed explicitly as follows
$f_{\theta}(x)=\frac{e^{-\gamma\theta}}{\Gamma(\theta)}\left(x^{\theta-1}+\sum_{k=1}^{[x]-1}\frac{(-\theta)^{k}}{k!}\idotsint\limits_{\begin{subarray}{c}t_{1},\ldots,t_{k}>1\\ t_{1}+\cdots+t_{k}\leq x\end{subarray}}(x-t_{1}-\cdots-t_{k})^{\theta-1}\frac{dt_{1}\ldots dt_{k}}{t_{1}\cdot\ldots\cdot t_{k}}\right),$
where $x>0$, $\cdot$] is the integer part function; see [GRI88, Thm. 1] and [GKL+24, Prop. 1] for similar representations. Moreover, it was shown in [CSZ19] that the following recurrent relation holds
$f_{\theta}(x)=\begin{cases}\frac{e^{-\gamma\theta}}{\Gamma(\theta)}x^{\theta-1},&0<x\leq 1,\\ \frac{e^{-\gamma\theta}}{\Gamma(\theta)}x^{\theta-1}-\theta x^{\theta-1}\int_{0}^{x-1}\frac{f_{\theta}(z)}{(1+z)^{\theta}}dz,&x>1.\end{cases}$
Note also that it follows from (1) that $X_{D}$ can be represented as
$X_{D}\overset{d}{=}U_{1}^{1/\theta}+(U_{1}U_{2})^{1/\theta}+(U_{1}U_{2}U_{3})^{1/\theta}+(U_{1}U_{2}U_{3}U_{4})^{1/\theta}+\cdots,$ (3)
where $\{U_{k}\}_{k=1}^{\infty}$ is the sequence of mutually independent identically distributed random variables with uniform distribution on $[0,1]$, see e.g. [VER79]. The series of the form (3) is commonly referred to as a perpetuity and is usually related to a random difference equation, see the equation (7) below. The random difference equations, which allow modelling discrete processes with a random discounting coefficient and random input at each time step, have been studied in a series of works, see, for example, [KES73, VER79, EG94].
The characteristic function of $X_{D}$ distributed according to $GD(\theta)$ is given by
$\psi_{D}(z)=\mathbb{E}e^{izX_{D}}=\exp\left\{\theta\int\limits_{0}^{1}(e^{izu}-1)\frac{du}{u}\right\},\penalty 10000\ z\in\mathbb{R},$ (4)
and it follows from (4) that the distribution $GD_{\theta}$ is infinitely divisible with Lévy measure
$\upsilon(du)=\frac{\theta}{u}\mathds{1}_{(0,1)}(u)du.$
The $GD_{\theta}$ distribution is also self-decomposable with the canonical function $k(x)$ given by $k(x)=\theta\mathds{1}_{(0,1)}(x)$ [KEN99, Cor. 15.11].
3 Dickman distributions over $\mathbb{R}^{d}$
The multidimensional Dickman distribution (as in Definition 3.1 below) was first introduced in [BM20] and further studied in [GZ24b], see also [GZ24a]. In this section, we generalise the definition given there to a new class of distributions, which we will refer to as the operator Dickman distributions. We further study its properties. In particular, we derive the expression for the characteristic function of the operator Dickman distributions and consequently show that they constitute a subset of infinitely divisible distributions over $\mathbb{R}^{d}$. We also demonstrate that the class of operator Dickman distributions forms a subset of the set of operator selfdecomposable distributions. Some specific examples of the distributions arising from this class are considered at the end of the section.
First, we introduce some notation which will be used throughout the paper. For elements $\mathbf{x}$ and $\mathbf{y}$ of the Euclidean space $\mathbb{R}^{d}$, $d\geq 1$, we denote their inner product by $\mathbf{x}\cdot\mathbf{y}$ and the Euclidean norm of $\mathbf{x}\in\mathbb{R}^{d}$ as $|\mathbf{x}|$. The set $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}:|\mathbf{x}|=1\}$ is the unit sphere in $\mathbb{R}^{d}$. For any topological space $S$ let $\mathscr{B}(S)$ denote the Borel sigma-algebra over $S$. Let also $\mathbb{R}_{+}$ denote the set of positive real numbers, and $\mathbb{C}$ denote the set of complex numbers. We will also use $\delta$ and $\mathds{1}$ as standard notation for the Dirac delta function and the indicator function, respectively; $a\wedge b$ denotes $\min\{a,b\}$ for $a,b\in\mathbb{R}$.
The linear transformations of $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ are usually identified with their matrix representations, and the set of all $d\times d$ matrices is denoted by $\mathscr{M}$. Additionally, $\mathscr{M_{+}}\subset\mathscr{M}$ ($\mathscr{M}_{-}\subset\mathscr{M}$) denotes the set of all elements of $\mathscr{M}$ whose eigenvalues have positive (negative) real parts. Let $I\in\mathscr{M}$ and $O\in\mathscr{M}$ denote, respectively, the identity and zero transformations, let also $\mathbb{I}=\{cI,c>0\}$ be a set of positive multiples of the identity transformation. For any transformation $Q\in\mathscr{M}$, let $Q^{*}$ denote its adjoint. For any $Q\in\mathscr{M}$, $t\in\mathbb{R}$ and $u\in\mathbb{R}_{+}$ define the matrix exponential as follows
$e^{tQ}\coloneq\sum_{k=0}^{\infty}\frac{t^{k}}{k!}Q^{k}\quad\text{and}\quad u^{Q}\coloneq e^{Q\log u}.$
Let $\|Q\|\coloneq\sup_{|\mathbf{x}|\leq 1}|Q\mathbf{x}|$ denote the operator norm of $Q\in\mathscr{M}$. For any measure $\vartheta$ on $\mathbb{R}^{d}$ and any invertible operator $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ denote $T\vartheta$ the measure given by $T\vartheta(A)\coloneq\vartheta(T^{-1}(A))$, $A\in\mathscr{B}(\mathbb{R}^{d})$.
Let $\mathscr{P}(\mathbb{R}^{d})$ denote the set of probability measures over $\mathbb{R}^{d}$ and for any $\pi\in\mathscr{P}(\mathbb{R}^{d})$ let $\hat{\pi}$ denote its characteristic function, i.e. $\hat{\pi}(\mathbf{z})=\int\nolimits_{\mathbb{R}^{d}}e^{i\mathbf{z}\cdot\mathbf{x}}\pi(d\mathbf{x})$. Let $H\subset\mathscr{P}(\mathbb{R}^{d})$ be the set of all probability measures $\nu$ such that
$\int\nolimits_{|\mathbf{x}|>1}\log|\mathbf{x}|\nu(d\mathbf{x})<\infty,$
and let $\mathscr{S}\subset\mathscr{P}(\mathbb{R}^{d})$ be the set of all probability distributions concentrated on $d$-dimensional sphere, i.e. $\sigma\in\mathscr{S}$ whenever $\sigma(\mathbb{S}^{d-1})=1$. Note that $\mathscr{S}\subset H$. For a random vector $\mathbf{X}\in\mathbb{R}^{d}$ let $\mathscr{L}(\mathbf{X})\in\mathscr{P}(\mathbb{R}^{d})$ denote its probability distribution.
We also denote $ID(\mathbb{R}^{d})\subset\mathscr{P}(\mathbb{R}^{d})$ the set of all infinitely divisible distributions. Recall that for any $\varrho\in ID(\mathbb{R}^{d})$ the characteristic function of $\varrho$ is given by the Lévy-Khintchine formula
$\hat{\varrho}(\mathbf{z})=\exp\left[i\mathbf{a}\cdot\mathbf{z}-\frac{1}{2}\mathbf{z}\cdot A\mathbf{z}+\int\limits_{\mathbb{R}^{d}}(e^{i\mathbf{z}\cdot\mathbf{x}}-1-i\mathbf{z}\cdot\mathbf{x}\mathds{1}_{|\mathbf{x}|\leq 1})\ell(d\mathbf{x})\right],\;\mathbf{z}\in\mathbb{R}^{d},$
where $\mathbf{a}\in\mathbb{R}^{d}$, $A\in\mathscr{M}$ is a symmetric non-negative definite matrix and $\ell$ is the Lévy measure on $\mathbb{R}^{d}$ satisfying $\ell(\{0\})=0$ and $\int_{\mathbb{R}^{d}}(1\wedge|\mathbf{x}|^{2})\ell(d\mathbf{x})<\infty$; the triplet $(\mathbf{a},A,\ell)$ is the characteristic triplet of the distribution $\varrho$, see [KEN99] for more details.
Let us present the definition of the multidimensional Dickman distribution, which was introduced in [BM20].
Definition 3.1 (Multidimensional Dickman distribution).
We call the distribution $\mathbb{D}_{\theta,\sigma}=\mathscr{L}(\mathbf{D})$ of the random vector $\mathbf{D}=\mathbf{D}_{\theta}^{(\sigma)}$ a multidimensional Dickman distribution if $\mathbf{D}$ satisfies the distributional equation
$\mathbf{D}\overset{d}{=}U^{1/\theta}(\mathbf{D}^{\prime}+\mathbf{V}),$ (5)
where $\theta>0$, $\mathbf{D}\overset{d}{=}\mathbf{D}^{\prime}$, $U$ has the uniform distribution on $[0,1]$, $\mathscr{L}(\mathbf{V})=\sigma\in\mathscr{S}$, and the random elements $U,\mathbf{V},\mathbf{D}^{\prime}$ are mutually independent.
Next, we generalise this definition to a wider class of distributions by substituting the scalar function $u^{1/\theta},\theta>0$ with a matrix exponential $u^{Q},Q\in\mathscr{M}_{+}$ and a probability measure $\sigma\in\mathscr{S}$ with a probability measure $\nu\in H$.
Definition 3.2.
For any $Q\in\mathscr{M}_{+}$ and $\nu\in H$, the probability distribution $\mathscr{D}(Q,\nu)\in\mathscr{P}(\mathbb{R}^{d})$ is the distribution of a vector $\mathbf{X}$, which satisfies
$\mathbf{X}\overset{d}{=}U^{Q}(\mathbf{X}^{\prime}+\mathbf{W}),$ (6)
where $\mathbf{X}\overset{d}{=}\mathbf{X}^{\prime}$, $U$ is uniformly distributed over $[0,1]$, $\mathscr{L}(\mathbf{W})=\nu$, and the random elements $U,\mathbf{X}^{\prime}$ and $\mathbf{W}$ are mutually independent.
Moreover, for any $\mathscr{Q}\subseteq\mathscr{M}_{+}$ and $\mathscr{H}\subseteq H$ let $\mathscr{D}(\mathscr{Q},\mathscr{H})$ be a set of distributions defined as
$\mathscr{D}(\mathscr{Q},\mathscr{H})=\{\mathscr{D}(Q,\nu),Q\in\mathscr{Q},\nu\in\mathscr{H}\},$
and the set $\mathscr{D}(\mathscr{\mathscr{M}}_{+},H)$ will be referred to as the class of operator Dickman distributions.
Remark 1.
Note that the random operator $U^{Q}$ is well-defined with probability one since the operator $s^{Q}$, $Q\in\mathscr{M}$, is well-defined for any $s>0$. The properties of the exponential operators, which we will use later, are investigated in [MS01, Ch. 2].
Notice that for any $Q\in\mathscr{M}_{+}$ the map $\gamma(a,\mathbf{b}):\cdot\mapsto a^{Q}(\cdot+\mathbf{b})$ with $a\in\mathbb{R}_{+}$ and $\mathbf{b}\in\mathbb{R}^{d}$ defines an affine transformation on $\mathbb{R}^{d}$. Moreover the family $\gamma=\{\gamma(a,\mathbf{b}),a>0,\mathbf{b}\in\mathbb{R}^{d}\}$ forms a non-commutative group under the composition, which is isomorphic to the group $\mathbb{R}_{+}\times\mathbb{R}^{d}$ with group operation * defined as $(a_{1},\mathbf{b}_{1})*(a_{2},\mathbf{b}_{2}):=(a_{1}a_{2},(a_{1}a_{2})^{Q}\mathbf{b}_{1}+a_{2}^{Q}\mathbf{b}_{2})$. We can now equip the group $\gamma$ with a probability measure defined as the product measure $Leb_{1}\times\nu$, where $Leb_{1}$ is the Lebesgue measure on $[0,1]$ and $\nu\in H$ is some probability measure. Therefore, the vector $\mathbf{X}$, defined in (6), can be viewed as the distributional fixed point of the random affine transformation described above. For $d=1$ this problem was studied in multiple works, see [VER79] for a short overview and references, and [BG19] for a generalisation to non-affine transformations.
As another point of view on Definition 3.2 we can consider the difference equation
$\mathbf{Y}_{n}=M_{n}\mathbf{Y}_{n-1}+\bm{\xi}_{n},\;n\geq 1,$ (7)
with some initial condition $\mathbf{Y}_{0}$ and random coefficients $M_{n}\in\mathscr{M}$ and $\bm{\xi}_{n}\in\mathbb{R}^{d}$, $n\geq 1$. Indeed, by iteration, it is easy to see that the solution of (6) can be rewritten as
$\mathbf{X}\overset{d}{=}\sum_{k=1}^{\infty}(U_{1}\cdot\ldots\cdot U_{k})^{Q}\mathbf{W}_{k},$ (8)
where $U_{1},U_{2},\ldots$ and $\mathbf{W}_{1},\mathbf{W}_{2},\dots$ are independent sequences of mutually independent random elements, sampled from the uniform distribution on $[0,1]$ and the measure $\nu$, respectively. Thus, the solution $\mathbf{X}$ of (6) is a weak limit as $n\to\infty$ of $\mathbf{Y}_{n}$ given by (7) in the case of $M_{k}=U_{k}^{Q}$ and $\bm{\xi}_{k}=M_{k}\mathbf{W}_{k}$.
The existence of a weak limit of $\mathbf{Y}_{n}$ as $n\to\infty$, independent of the initial condition $\mathbf{Y}_{0}$, under some mild conditions on $M_{n}$ and $\bm{\xi}_{n}$ was established in the celebrated work of Kesten [KES73]. In particular, Kesten [KES73, Thm. 6] showed that the limit exists if $\infty<\mathbb{E}\log\|M_{1}\|<0$ and $0<\mathbb{E}|\bm{\xi}_{1}|^{\kappa}<\infty$ for some $\kappa>0$. However, as shown in [VER72] for $d=1$, the last condition can be weakened to $\mathbb{E}\log^{+}|\xi_{1}|<\infty$ under the assumption that the first condition holds, and this proof translates easily for $d>1$. Hereafter, we define
$\log^{+}x=\begin{cases}0,&0<x\leq 1\\ \log x,&x>1.\end{cases}$
The model (7) is also commonly referred to as the random-coefficient $AR(1)$ process and was proposed as a valid model in physical, biological and financial applications, see [VER79] and references therein. The one-dimensional version of this problem has been studied in a series of works, see e.g. [CS75, GRI75, EG94].
Another interpretation of the representation (8) can be obtained if we rewrite first
$\mathbf{X}\overset{d}{=}\sum_{k=1}^{\infty}f(T_{k})\mathbf{W}_{k},$ (9)
where $T_{k}$ are the arrival times of the Poisson process with intensity $1$ and $f:\mathbb{R}_{+}\mapsto\mathscr{M}$ is the measurable function defined by $f(t)=e^{-tQ}$. The sum in (9) represents the weak limit of the sum
$\sum_{k=1}^{\infty}f(T_{k})\mathds{1}\{T_{k}<t\}\mathbf{W}_{k}$ (10)
as $t\to\infty$. The equation (10) can be interpreted as the accumulation, up to time $t$, of a multidimensional signal. This signal appears at random time points according to the Poisson process with intensity $1$; the amplitudes are independent of each other and of the arrival times, and identically distributed according to the law $\nu$; the signal propagates in time according to the law $f$. For more details on such models, we refer to [TAK54, TAK55, WES76].
Remark 2.
It is evident that the set of multidimensional Dickman distributions $\{\mathbb{D}_{\theta,\sigma},\theta>0,\sigma\in\mathscr{S}\}$ coincides with a subset $\mathscr{D}(\mathbb{I},\mathscr{S})$ of $\mathscr{D}(\mathscr{\mathscr{M}}_{+},H)$. Moreover, for $d=1$ the sphere $\mathbb{S}^{0}$ consists of two points $\mathbb{S}^{0}=\{-1,1\}$ and by choosing in this case $Q=\theta>0$ and $\sigma(\{1\})=1-\sigma(\{-1\})=1$ we can recover the Dickman distribution $GD_{\theta}$.
The one-dimensional distribution $\mathscr{D}(1/\theta,\nu)$ with a more general type of probability measure $\nu\in H$ appears naturally in the contexts of shot-noise processes and random difference equations as well as in the limit schemes, see, e.g, [CS75, VER79, PIN18]. Notably, in the particular case $\nu(dx)=e^{-x}dx,x>0$, the distribution $\mathscr{D}(1/\theta,\nu)$ coincides with the well-known Gamma distribution having characteristic function of the form
$\exp\left\{\theta\int\limits_{0}^{1}(e^{izu}-1)e^{-u}\frac{du}{u}\right\},\penalty 10000\ z\in\mathbb{R}.$
Remark 3.
Note that in the case of $\nu=\delta_{\mathbf{0}}$ it holds that $\mathbf{X}=\mathbf{0}$ almost surely, and this situation is trivial. We shall exclude it from consideration and, furthermore, assume that $\nu(\{\mathbf{0}\})=0$.
To establish properties of the operator Dickman distributions, we will need the following lemma.
Lemma 3.3.
Let $Q\in\mathscr{M}_{+}$ be fixed.
  1. (i)
    For all $\mathbf{x}\in\mathbb{R}^{d}$ the integral
    $P_{1}(\mathbf{x})=\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|\wedge 1)\frac{ds}{s}$
    is finite, and $P_{1}(\mathbf{x})\leq B_{1}\log(1+|\mathbf{x}|)$ for some $B_{1}>0$.
  2. (ii)
    For all $\mathbf{x}\in\mathbb{R}^{d}$ the integral
    $P_{2}(\mathbf{x})=\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|^{2}\wedge 1)\frac{ds}{s}$
    is finite, and $P_{2}(\mathbf{x})\leq B_{2}\log(1+|\mathbf{x}|^{2})$ for some $B_{2}>0$.
Proof.
(i) First note that the function $f(t)=t\wedge 1$ is non-decreasing. Note also that for any $Q\in\mathscr{M}_{+}$ there exist $K>0$ and $c_{1}>0$ such that $|s^{Q}\mathbf{x}|\leq c_{1}s^{K}|\mathbf{x}|$ for all $s>0$ and $\mathbf{x}\in\mathbb{R}^{d}$ [MS01, Prop. 2.2.11], see also the proof of [URB72, Lem. 6.1.]. We might as well choose $c_{1}>1$, thus
$\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|\wedge 1)\frac{ds}{s}$ $\leq c_{1}\int\limits_{0}^{1}(s^{K}|\mathbf{x}|\wedge 1)\frac{ds}{s}=\frac{c_{1}}{K}\int\limits_{0}^{1}(s|\mathbf{x}|\wedge 1)\frac{ds}{s}$
$=\frac{c_{1}}{K}\int\limits_{0}^{|\mathbf{x}|}(s\wedge 1)\frac{ds}{s}=\frac{c_{1}}{K}\left(1\wedge|\mathbf{x}|-\log\left(1\wedge|\mathbf{x}|^{-1}\right)\right),$
and the result follows.
(ii) This is a direct consequence of [URB72, Lem. 6.1]
Proposition 3.4.
For any $Q\in\mathscr{M}_{+}$ and $\nu\in H$, the distribution $\mathscr{D}=\mathscr{D}(Q,\nu)$ has the characteristic function of the form
$\psi(\mathbf{z})=\hat{\mathscr{D}}(\mathbf{z})=\exp\left[\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}\right],\;\mathbf{z}\in\mathbb{R}^{d}.$ (11)
Proof.
Note first that the finiteness of the integral in (11) follows by Lemma 3.3i and the assumption $\nu\in H$. Notice now that (6) yields
$\psi(\mathbf{z})=\int\limits_{0}^{1}\psi(u^{Q^{*}}\mathbf{z})\hat{\nu}(u^{Q^{*}}\mathbf{z})du.$
Replacing the variable $\mathbf{z}$ with $r^{Q^{*}}\bm{\omega}$ for $r\geq 0$ and $\bm{\omega}\in\mathbb{R}^{d}$, where by convention $0^{Q^{*}}\bm{\omega}=\mathbf{0}$ for all $\bm{\omega}\in\mathbb{R}^{d}$, we get
$\psi(r^{Q^{*}}\bm{\omega})=$ $\int\limits_{0}^{1}\psi((ur)^{Q^{*}}\bm{\omega})\hat{\nu}((ur)^{Q^{*}}\bm{\omega})du$
$=$ $\frac{1}{r}\int\limits_{0}^{r}\psi(u^{Q^{*}}\bm{\omega})\hat{\nu}(u^{Q^{*}}\bm{\omega})du.$
Since the latter expression is differentiable with respect to $r$, we can obtain
$\frac{\partial}{\partial r}\psi(r^{Q^{*}}\bm{\omega})=$ $-\frac{1}{r^{2}}\int\limits_{0}^{r}\psi(u^{Q^{*}}\bm{\omega})\hat{\nu}(u^{Q^{*}}\bm{\omega})du+\frac{1}{r}\psi(r^{Q^{*}}\bm{\omega})\hat{\nu}(r^{Q^{*}}\bm{\omega})$
$=$ $\frac{1}{r}\psi(r^{Q^{*}}\bm{\omega})(\hat{\nu}(r^{Q^{*}}\bm{\omega})-1),$
or, alternatively,
$\frac{\partial}{\partial r}\log\psi(r^{Q^{*}}\bm{\omega})=\frac{1}{r}(\hat{\nu}(r^{Q^{*}}\bm{\omega})-1).$
It follows from the above that there exists $g:\mathbb{R}^{d}\mapsto\mathbb{C}$ such that
$\log\psi(r^{Q^{*}}\bm{\omega})=\int\limits_{0}^{r}(\hat{\nu}(s^{Q^{*}}\bm{\omega})-1)\frac{ds}{s}+g(\bm{\omega})=\int\limits_{0}^{1}(\hat{\nu}((sr)^{Q^{*}}\bm{\omega})-1)\frac{ds}{s}+g(\bm{\omega}).$
Letting $r=0$ we can deduce from $\psi(\mathbf{0})=1$ that $g(\bm{\omega})=0$ for all $\bm{\omega}\in\mathbb{R}^{d}$.
Returning to the variable $\mathbf{z}$, we finally get
$\log\psi(\mathbf{z})=\int\limits_{0}^{1}(\hat{\nu}(s^{Q^{*}}\mathbf{z})-1)\frac{ds}{s}=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{is^{Q^{*}}\mathbf{z}\cdot\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s},$
and the result follows since $\left(s^{Q^{*}}\right)^{*}=s^{Q}$. ∎
Corollary 3.5.
The characteristic function $\psi(\mathbf{z})$ of $\mathscr{D}(Q,\nu)$, $Q\in\mathscr{M}_{+}$, $\nu\in H$, can be written as
$\psi(\mathbf{z})=\exp\left[\int\limits_{0}^{1}(\hat{\nu}(s^{Q^{*}}\mathbf{z})-1)\frac{ds}{s}\right].$
Corollary 3.6.
The class $\mathscr{D}(\mathscr{M}_{+},H)$ is a subset of $ID(\mathbb{R}^{d})$ and for any $Q\in\mathscr{M}_{+}$ and $\nu\in H$ the distribution $\mathscr{D}(Q,\nu)$ has the characteristic triplet $(\mathbf{a},O,M)$, where
$\begin{gathered}\mathbf{a}=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}s^{Q}\mathbf{x}\mathds{1}_{|s^{Q}\mathbf{x}|\leq 1}\nu(d\mathbf{x})\frac{ds}{s},\\ M(B)=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}\mathds{1}_{B}(s^{Q}\mathbf{x})\nu(d\mathbf{x})\frac{ds}{s},\;B\in\mathscr{B}(\mathbb{R}^{d}).\end{gathered}$ (12)
Proof.
The result follows directly from (11) after verification that $|\mathbf{a}|<\infty$ and $\int_{\mathbb{R}^{d}}(1\wedge|\mathbf{y}|^{2})M(d\mathbf{y})<\infty$. Indeed, using Lemma 3.3i we obtain that
$|\mathbf{a}|\leq\int\limits_{0}^{1}\int\limits_{\mathbb{R}^{d}}(|s^{Q}\mathbf{x}|\wedge 1)\nu(d\mathbf{x})\frac{ds}{s}\leq B_{1}\int\limits_{\mathbb{R}^{d}}\log(1+|\mathbf{x}|)\nu(d\mathbf{x})<\infty.$
Similarly, using Lemma 3.3ii, we get
$\int\limits_{\mathbb{R}^{d}}(1\wedge|\mathbf{y}|^{2})M(d\mathbf{y})=$ $\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(1\wedge|s^{Q}\mathbf{x}|^{2})\nu(d\mathbf{x})\frac{ds}{s}$
$\leq$ $B_{2}\int\limits_{\mathbb{R}^{d}}\log(1+|\mathbf{x}|^{2})\nu(d\mathbf{x})<\infty.$
Note that we used the assumption $\nu\in H$ in both parts. ∎
Remark 4.
Remarkably, the distributions with Lévy measure as in (12) have previously appeared in [JUR82].
Proposition 3.7.
  1. (i)
    Assume that $\nu\in H$ is such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(\mathbf{x})<\infty$. Then the mean vector $\mathbf{m}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and is given as
    $\mathbf{m}_{Q,\nu}=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}s^{Q}\mathbf{x}\nu(d\mathbf{x})\frac{ds}{s}.$ (13)
  2. (ii)
    Assume that $\nu\in H$ such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(\mathbf{x})<\infty$. Then the covariance matrix $\mathbb{C}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and can be given by
    $\mathbb{C}_{Q,\nu}=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(s^{Q}\mathbf{x})(s^{Q}\mathbf{x})^{T}\nu(d\mathbf{x})\frac{ds}{s},$ (14)
    where $\mathbf{x}\mathbf{y}^{T}=[x_{i}y_{j}]_{i,j=1}^{d}\in\mathscr{M}$ for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$.
Proof.
Both formulas (13) and (14) follow easily by differentiation of the characteristic function (11). Note that the finiteness of the integrals in (13) under the assumptions of (i) follows again by the estimate
$|s^{Q}\mathbf{x}|\leq{c_{1}}s^{K}|\mathbf{x}|,\;s>0,\mathbf{x}\in\mathbb{R}^{d}$
with some $c_{1},K>0$. Consequently, we obtain
$|\mathbf{m}|\leq\frac{c_{1}}{K}\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty.$
The finiteness of the integrals in (14) follows similarly. ∎
Remark 5.
Note that it follows from
$\frac{d}{dt}e^{tX}=Xe^{tX}=e^{tX}X,\quad t\in\mathbb{R},\penalty 10000\ X\in\mathscr{M},$
and integration by parts, that formulas (13) and (14) can be rewritten as
$\begin{gathered}\mathbf{m}_{Q,\nu}=Q^{-1}\int\limits_{\mathbb{R}^{d}}\mathbf{x}\nu(d\mathbf{x})\;\text{ and }\;\mathbb{C}_{Q,\nu}=\mathscr{A}_{Q}^{-1}\left(\int\limits_{\mathbb{R}^{d}}\mathbf{x}\mathbf{x}^{T}\nu(d\mathbf{x})\right),\end{gathered}$
where $\mathscr{A}_{Q}:\mathscr{M}_{d}\mapsto\mathscr{M}_{d}$, $\mathscr{A}_{Q}(B)=QB+BQ^{*}$.
Proposition 3.8.
  1. (i)
    For any fixed $Q\in\mathscr{M}_{+}$, the class $\mathscr{D}(Q,H)$ is closed under transformations $\mathbf{x}\mapsto t^{A}\mathbf{x}$ for any $t>0$ and $A\in\mathscr{M}$ commuting with Q (i.e. $AQ=QA$). In particular, $\mathscr{D}(Q,H)$ is closed under transformations $\mathbf{x}\mapsto t^{\alpha}\mathbf{x}$ for any $t>0$ and $\alpha>0$.
  2. (ii)
    Consider the distribution $\mathscr{D}(\frac{1}{\theta}I,\nu)$, $\theta>0$, $\nu\in H$ and assume that the distribution $\nu$ is invariant under the action of the operator $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, i.e. $T\nu=\nu$. Then $\mathscr{D}(\frac{1}{\theta}I,\nu)$ is also invariant under $T$. In particular, if $\nu$ is rotationally invariant, $\mathscr{D}(\frac{1}{\theta}I,\nu)$ is also rotationally invariant for any $\theta>0$.
  3. (iii)
    For any fixed $Q\in\mathscr{M}_{+}$ the class $\mathscr{D}(\frac{1}{c}Q,H)$, where $c$ runs over $(0,\infty)$, is closed under finite convolutions.
Proof.
(i) Let $Q\in\mathscr{M}_{+}$ be fixed and let $\mathbf{X}$ be a random vector such that $\mathscr{L}(\mathbf{X})\in\mathscr{D}(Q,H)$. Then there exists a random vector $\mathbf{W}$, $\mathscr{L}(\mathbf{W})\in H$, such that (6) holds. Therefore, $t^{A}\mathbf{X}$ satisfies
$t^{A}\mathbf{X}\overset{d}{=}t^{A}U^{Q}(\mathbf{X}+\mathbf{W})$
for any $t>0$. Since $A$ and $Q$ commute, then $s^{A}$ and $t^{Q}$ also commute for any $s,t>0$. Thus, we obtain
$t^{A}\mathbf{X}\overset{d}{=}U^{Q}(t^{A}\mathbf{X}+t^{A}\mathbf{W}),$
and, moreover, $\mathbb{E}\log^{+}|t^{A}\mathbf{W}|\leq\mathbb{E}\log^{+}(\|t^{A}\|\cdot|\mathbf{W}|)<\infty$. This implies that $\mathscr{L}(t^{A}\mathbf{X})\in\mathscr{D}(Q,H)$ or, more specifically, $t^{A}\mathscr{D}(Q,\nu)=\mathscr{D}(Q,t^{A}\nu)$.
The proof of part (ii) is analogous.
(iii) Note first that for fixed $Q\in\mathscr{M}_{+}$, $c>0$ and $\mathscr{D}=\mathscr{D}(\frac{1}{c}Q,\nu)$, from (11) by a change of variables we obtain
$\log\widehat{\mathscr{D}}(\mathbf{z})=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{\frac{1}{c}Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}=c\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}.$
Let $Q\in\mathscr{M}_{+}$, $c_{1},c_{2}>0$ and $\nu_{1},\nu_{2}\in H$ and consider the independent distributions $\mathscr{D}_{1}=\mathscr{D}(\frac{1}{c_{1}}Q,\nu_{1})$, $\mathscr{D}_{2}=\mathscr{D}(\frac{1}{c_{2}}Q,\nu_{2})$. If we denote $\mathscr{D}=\mathscr{D}_{1}*\mathscr{D}_{2}$, then
$\log\widehat{\mathscr{D}}(\mathbf{z})$ $=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)(c_{1}\nu_{1}+c_{2}\nu_{2})(d\mathbf{x})\frac{ds}{s}$
$=(c_{1}+c_{2})\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\frac{(c_{1}\nu_{1}+c_{2}\nu_{2})(d\mathbf{x})}{c_{1}+c_{2}}\frac{ds}{s},$
and it is easy to see that $\mathscr{D}=\mathscr{D}(\frac{1}{c_{1}+c_{2}}Q,\frac{c_{1}\nu_{1}+c_{2}\nu_{2}}{c_{1}+c_{2}})$. ∎
Remark 6.
The case of countable convolutions is considered in Corollary 4.6.
Let us now show that each $\mathscr{D}\in\mathscr{D}(Q,H)$ is $Q$-selfdecomposable. The definition of operator selfdecomposability was first given in [URB72]. Following [URB72], we say that a measure (distribution) on $\mathbb{R}^{d}$ is full if its support is not contained in any (d1)(d-1)-dimensional hyperplane. Urbanik proved that each full $Q$-selfdecomposable distribution appears as a weak limit of some triangular array $A_{n}\sum_{k=1}^{n}\mathbf{X}_{k}+\mathbf{a}_{n},$ $n\in\mathbb{N}$, where $\mathbf{X}_{n},$ $n\geq 1$ is a sequence of independent $\mathbb{R}^{d}$-valued random variables, $A_{n}\in\mathscr{M}_{-}$ and $\mathbf{a}_{n}\in\mathbb{R}^{d},$ $n\geq 1$. The class of operator selfdecomposable distributions was studied in [URB72, SY84, MAS04], see also [SY85, MS01] for the subclass of operator-stable distributions and applications in limit theorems.
Let us recall the definition of operator selfdecomposable distributions.
Definition 3.9 ([URB72]).
The distribution $\varrho\in\mathscr{P}(\mathbb{R}^{d})$ is $Q$-selfdecomposable for some $Q\in\mathscr{M}_{+}$ if for every $t\in\mathbb{R}_{+}$ there exists $\varrho_{t}\in\mathscr{P}(\mathbb{R}^{d})$ such that
$\hat{\varrho}(\mathbf{z})=\hat{\varrho}(e^{-tQ^{*}}\mathbf{z})\hat{\varrho}_{t}(\mathbf{z}).$
In [SY84], the authors presented two useful representations for the characteristic function of the operator selfdecomposable distribution, which allow us to establish the following proposition.
Proposition 3.10.
Let $Q\in\mathscr{M}_{+}$ and $\nu\in H$ be given, then $\mathscr{D}=\mathscr{D}(cQ,\nu)$, is $Q$-selfdecomposable for any $c>0$.
Proof.
It follows easily from (11) and representations given in [SY84]. Indeed, for each $\mathscr{D}(Q,\nu)$, we can choose $\rho=\nu$ in [SY84, Second representation theorem of $Q$-selfdecomposable distributions]. ∎
Note that the classes of $Q$-selfdecomposable Dickman distributions $\mathscr{D}(Q,H)$ might have non-empty intersections even for distinct $Q\in\mathscr{M}_{+}$, as proved in the proposition below. This question for the case of operator stable distributions was studied in [HM81].
Proposition 3.11.
Consider $Q\in\mathscr{M}_{+}$ and assume that $Q$ has at least one real eigenvalue. Assume also $Q\notin\mathbb{I}$. Consider now $\mathscr{D}=\mathscr{D}(Q,\nu)$, with $\nu$ being the probability measure supported on the real eigenspaces of $Q$. Then also $\mathscr{D}\in\mathscr{D}(\mathbb{I},H)$ and, consequently, $\mathscr{D}$ is both $Q$- and $I$-selfdecomposable.
Proof.
Let $a_{1},\ldots,a_{n}>0$, $n\geq 1$, denote the distinct real eigenvalues of $Q$ and $A_{1},\ldots,A_{n}$ be the corresponding eigenspaces. Thus, we have $Q\mathbf{x}=a_{k}\mathbf{x}$ and $s^{Q}\mathbf{x}=s^{a_{k}}\mathbf{x}$ for any $\mathbf{x}\in A_{k}$, $k=1,\ldots,n$. Denote also $p_{k}=\nu(A_{k})$, $p_{k}\geq 0$, $k=1,\ldots,n$ and note that $p_{1}+\cdots+p_{n}=1$ by the assumptions above. It follows from Proposition 3.4 that
$\hat{\mathscr{D}}(\mathbf{z})=$ $\exp\left[\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}\right]$
$=$ $\exp\left[\sum_{k=1}^{n}\int\limits_{A_{k}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{a_{k}}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}\right]$
$=$ $\exp\left[\sum_{k=1}^{n}\int\limits_{A_{k}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s\mathbf{x}}-1)a_{k}^{-1}\nu(d\mathbf{x})\frac{ds}{s}\right].$
Let $\nu_{k}$ be the restriction of the measure $\nu$ to $A_{k}$, $k=1,\ldots,n$. Rewrite now
$\hat{\mathscr{D}}(\mathbf{z})=$ $\exp\left[\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s\mathbf{x}}-1)\sum_{k=1}^{n}a_{k}^{-1}\nu_{k}(d\mathbf{x})\frac{ds}{s}\right]$
$=$ $\exp\left[\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{\theta}\mathbf{x}}-1)\tilde{\nu}(d\mathbf{x})\frac{ds}{s}\right],$
[…]
Let us now consider some specific examples of the distributions in the class of operator Dickman distributions.
Example 3.12.
Let us fix $\Lambda\in\mathscr{M}_{+}$ and assume that $\Lambda$ is diagonal, that is $\Lambda=diag(a_{1},\ldots,a_{d})$ for $a_{i}>0,i=1,\ldots,d$. We will also assume that $\Lambda\notin\mathbb{I}$. Let $\mathscr{L}(\mathbf{X})=\mathscr{D}(\Lambda,\nu)$ for some $\nu\in H$. Since in this case it holds for $t>0$ that $t^{\Lambda}=diag(s^{a_{1}},\ldots,s^{a_{d}})$ it is clear that the components $X_{i}$ of the vector $\mathbf{X}$ satisfy the distributional equation
$X_{i}\overset{d}{=}U^{a_{i}}(X_{i}+W_{i}),\;i=1,\ldots,d,$
where $W_{i},$ $i=1,\ldots,d$ are components of the random element $\mathbf{W}$, $\mathscr{L}(\mathbf{W})=\nu$. The random variables of this type have been considered in [CS75, VER79, PIN18]. Moreover, the formula (14) for the covariance matrix in this case simplifies to
$C_{Q,\nu}]_{i,j}=\frac{1}{a_{i}+a_{j}}\int\limits_{\mathbb{R}^{d}}x_{i}x_{j}\nu(d\mathbf{x}),\;i,j=1\ldots,d$
whenever $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(\mathbf{x})<\infty$. Notice also that in the simplest case $\nu=\delta_{\mathbf{w}}$, for some $\mathbf{w}=(w_{1},\ldots,w_{d})\in\mathbb{R}^{d}_{+}$, the components $w_{i}X_{i}$ are distributed according to $GD_{1/a_{i}}$, $i=1,\ldots,d$.
Example 3.13.
Let us assume now that $Q\in\mathscr{M}_{+}$ can be represented as $Q=S\Lambda S^{-1}$ for $\Lambda$ as in Example 3.12 and some invertible matrix $S$. Then we have $t^{Q}=St^{\Lambda}S^{-1}$ for $t>0$. Let again $\mathscr{L}(\mathbf{X})=\mathscr{D}(Q,\nu)$, $\nu\in H$ and $\mathscr{L}(\mathbf{W})=\nu$. Then it follows from (6) that
$\mathbf{X}\overset{d}{=}SU^{\Lambda}S^{-1}(\mathbf{X}+\mathbf{W}),$
or, alternatively,
$S^{-1}\mathbf{X}\overset{d}{=}U^{\Lambda}(S^{-1}\mathbf{X}+S^{-1}\mathbf{W}),$
and, therefore, $\mathscr{L}({S^{-1}\mathbf{X}})=\mathscr{D}(\Lambda,S^{-1}\nu)$, cf. Example 3.12.
Example 3.14.
Consider $\mathscr{D}=\mathscr{D}(\frac{1}{\theta}I,\nu)$, $\theta>0$ and assume $\nu\in\mathscr{S}$. As was noted before, the random vector $\mathbf{X}$ with $\mathscr{L}(\mathbf{X})=\mathscr{D}$ has the multidimensional Dickman distribution as defined in [BM20]. The properties of this distribution have been studied in [GZ24b], see also [GZ24a]. In particular, it was established that if the measure $\nu$ is finitely supported, then $\mathbf{X}$ can be expressed using independent Dickman random variables. Indeed, assume that $\nu$ is of the form
$\nu=\sum\limits_{i=1}^{n}p_{i}\delta_{\mathbf{w}_{i}},$ (15)
where $\mathbf{w}_{1},\ldots,\mathbf{w}_{n}\in\mathbb{S}^{d-1}$ and $p_{1},\ldots,p_{n}>0,\;p_{1}+\cdots+p_{n}=1$, $n\geq 1$. Then $\mathbf{X}$ can be represented as
$\mathbf{X}\overset{d}{=}\sum\limits_{i=1}^{n}\mathbf{w}_{i}X_{D}^{(i)},$ (16)
where $X_{D}^{(1)},\ldots,X_{D}^{(n)}$ are independent random variables and $X_{D}^{(i)}$ is distributed according to $GD_{\theta p_{i}}$, $i=1,\ldots,n$. The representation (16) was obtained in [GZ24b]. Notice also that it also holds for $\nu\in H$. Moreover, the representation (16) can be used for simulation from the multivariate Dickman distribution $\mathscr{D}(\frac{1}{\theta}I,\nu)$ as discussed in [GZ24b].
[…]
In what follows, we focus on obtaining relations for the probability density function of the operator Dickman distributions. Notice first that if the distribution $\mathscr{D}(Q,\nu)$, $Q\in\mathscr{M}_{+}$, $\nu\in H$, over $\mathbb{R}^{d}$ is full, then it is absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{d}$ [YAM83]. We also refer to [KEN99, Prop. 28.1] for sufficient conditions ensuring differentiability of the density for the distributions in the set $\mathscr{D}(\mathbb{I},\nu)$. Nevertheless, an explicit expression for the density is generally difficult to obtain. In this work, the explicit formula for the density is obtained for a specific case of the multidimensional Dickman distribution $\mathscr{D}\in\mathscr{D}(\mathbb{I},\mathscr{S})$ in Theorem 3.15, whereas Proposition 3.16 treats a more general case of $\mathscr{D}(\mathbb{I},H)$. Note that we utilise some elements of the theory of generalised functions operating on the Schwartz space of rapidly decreasing functions; for more details on this topic, see the classic book [GS69] and also [MIT13].
We consider first the multidimensional Dickman distribution $\mathscr{D}(\frac{1}{\theta}I,\nu)$, where the distribution $\nu$ is uniform on the sphere $\mathbb{S}^{d-1}$. In the following theorem, we obtain its density function, which can be viewed as the multivariate analogue of the density $f_{\theta}(\cdot)$, see Section 2. The idea of the proof is due to [VER72, Thm. 4.7.7].
We further denote $J_{a}(\cdot),a>-\frac{1}{2}$ the Bessel function of first kind, $Leb$ the Lebesgue measure on $\mathbb{S}^{d-1}$, $\zeta_{d}=(Leb(\mathbb{S}^{d-1}))^{-1}={\Gamma(d/2)}\cdot(2\pi)^{-d/2}$ the reciprocal surface area of $\mathbb{S}^{d-1}$ and $Y_{d}(s)=(s/2)^{1-\frac{d}{2}}\Gamma(d/2)J_{d/2-1}(s)$, $s>0$, $d>1$, the spherical Bessel function, see e.g. [PIN93]. Moreover, for $d=1$, we let $Y_{1}(s)=\cos(s),s>0$.
Theorem 3.15.
Let $d\geq 1$, $\theta>0,\theta\neq d+2n,n=0,1,\ldots$, and consider the distribution $\mathscr{D}=\mathscr{D}(\frac{1}{\theta}I,\nu)$. Assume also that the measure $\nu$ is uniform on $\mathbb{S}^{d-1}$, i.e. $\nu(B)=\zeta_{d}Leb(B)$ for any $B\in\mathscr{B}(\mathbb{S}^{d-1})$. Let $\mathscr{L}(\mathbf{X})=\mathscr{D}$. Then the density $f(\mathbf{x})$, $\mathbf{x}\in\mathbb{R}^{d}$ of $\mathbf{X}$ exists and can be given by
$f(\mathbf{x})=e^{\theta\alpha_{d}}\left(g(\mathbf{x})+\sum_{n=1}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}\idotsint\limits_{D_{n}}g(\mathbf{x}-\mathbf{x}_{1}-\cdots-\mathbf{x}_{n})\frac{d\mathbf{x}_{1}\ldots d\mathbf{x}_{n}}{|\mathbf{x}_{1}|^{d}\ldots|\mathbf{x}_{n}|^{d}}\right),$ (17)
where
$\alpha_{d}=\int\limits_{0}^{1}\left(Y_{d}(s)-1\right)\frac{ds}{s}+\int\limits_{1}^{\infty}Y_{d}(s)\frac{ds}{s}<\infty,$ (18)
$g(\mathbf{y})=2^{-\theta}\pi^{-d/2}\frac{\Gamma(\frac{d-\theta}{2})}{\Gamma(\frac{\theta}{2})}|\mathbf{y}|^{\theta-d},\penalty 10000\ \mathbf{y}\in\mathbb{R}^{d},$
$D_{n}=\left\{(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})\in\mathbb{R}^{d\times n}:|\mathbf{x}_{i}|>1,\penalty 10000\ i=1,\ldots,n\right\}.$
Proof.
Let us consider the function $\varphi(\mathbf{z})=\log\hat{\mathscr{D}}(\mathbf{z})$. Using Corollary 3.5 we obtain
$\varphi(\mathbf{z})=\theta\int_{0}^{1}(\hat{\nu}(s\mathbf{z})-1)\frac{ds}{s}.$
Moreover, for $\nu=\zeta_{d}Leb(B)$ we have
$\hat{\nu}(\mathbf{w})=Y_{d}(|\mathbf{w}|),\;\mathbf{w}\in\mathbb{R}^{d},$
see, for example, [PIN93]. Therefore, we obtain
$\theta^{-1}\varphi(\mathbf{z})$ $=\int\limits_{0}^{1}\left(Y_{d}(s|\mathbf{z}|)-1\right)\frac{ds}{s}=\int\limits_{0}^{|\mathbf{z}|}\left(Y_{d}(s)-1\right)\frac{ds}{s}$
$=\int\limits_{0}^{1}\left(Y_{d}(s)-1\right)\frac{ds}{s}+\int\limits_{1}^{\infty}Y_{d}(s)\frac{ds}{s}-\int\limits_{|\mathbf{z}|}^{\infty}Y_{d}(s)\frac{ds}{s}-\log|\mathbf{z}|.$
Since the function $Y_{d}(\cdot)$ satisfies
$Y_{d}(s)=1+O(s^{2}),\;s\to 0;$
$Y_{d}(s)=\left(\frac{s}{2}\right)^{\frac{1-d}{2}}\Gamma\left(\frac{d}{2}\right)\pi^{-1/2}\left[\cos\left(s-\frac{\pi(d-1)}{4}\right)+O\left(\frac{1}{s}\right)\right],\;s\to\infty,$
we get (18), see [PIN93, Prop. 2.1].
Therefore, we can rewrite
$\hat{\mathscr{D}}(\mathbf{z})=$ $\exp\left(\theta\alpha_{d}+\log|\mathbf{z}|^{-\theta}-\theta\int\limits_{1}^{\infty}Y_{d}(s|\mathbf{z}|)\frac{ds}{s}\right)$
$=$ $e^{\theta\alpha_{d}}|\mathbf{z}|^{-\theta}\exp\left(-\theta\int\limits_{1}^{\infty}Y_{d}(s|\mathbf{z}|)\frac{ds}{s}\right).$
Denote now $\tilde{F}_{k}(s)=s^{-d}\mathds{1}\{1<s<k\}$, $k\geq 2$ and $\tilde{F}(s)=s^{-d}\mathds{1}\{s>1\}$. Let also $F_{k}(\mathbf{x})=\tilde{F}_{k}(|\mathbf{x}|)$ and $F(\mathbf{x})=\tilde{F}(|\mathbf{x}|)$. Note that $\{F_{k},k\geq 2\}$ is a sequence of functions such that
  1. 1.
    $F_{k}(\mathbf{x})\in L_{1}(\mathbb{R}^{d})\cap L_{2}(\mathbb{R}^{d})$,
  2. 2.
    $F_{k}(\mathbf{x})\overset{L_{2}}{\to}F(\mathbf{x})$ as $k\to\infty$,
  3. 3.
    $F_{k}$ is radial for any $k\geq 2$ and
    $\int_{\mathbb{R}^{d}}e^{i\mathbf{z}\cdot\mathbf{x}}F_{k}(\mathbf{x})d\mathbf{x}=(2\pi)^{d/2}|\mathbf{z}|^{1-d/2}\int_{0}^{\infty}\tilde{F}_{k}(s)J_{d/2-1}(s|\mathbf{z}|)s^{d/2}ds,$
    see [SW71]; note that for $d=1$ this relation is obvious due to properties of even functions.
Thus, the function $\theta\int\nolimits_{1}^{\infty}Y_{d}(s|\mathbf{z}|)\frac{ds}{s}$ is the Fourier transform of the function $\theta\zeta_{d}F(\mathbf{x})\in L_{2}(\mathbb{R}^{d})$ (see [SW71] for more details on the theory of the Fourier transform on $L_{2}$), and we can write
$\theta\int\limits_{1}^{\infty}Y_{d}(s|\mathbf{z}|)\frac{ds}{s}=-\theta\zeta_{d}\int\limits_{\mathbb{R}^{d}}e^{i\mathbf{z}\cdot\mathbf{x}}|\mathbf{x}|^{-d}\mathds{1}\{|\mathbf{x}|>1\}d\mathbf{x},$
understanding the integral on the right as a limit in $L_{2}$ of integrals
$\int_{\mathbb{R}^{d}}e^{i\mathbf{z}\cdot\mathbf{x}}|\mathbf{x}|^{-d}\mathds{1}\{1<|\mathbf{x}|<k\}d\mathbf{x},\;k\to\infty.$
We arrive at
$e^{-\theta\alpha_{d}}\hat{\mathscr{D}}(\mathbf{z})=$ $|\mathbf{z}|^{-\theta}\exp\left(-\theta\zeta_{d}\int\limits_{|\mathbf{x}|>1}e^{i\mathbf{z}\cdot\mathbf{x}}|\mathbf{x}|^{-d}d\mathbf{x}\right)$
$=$ $|\mathbf{z}|^{-\theta}\left(1+\sum_{n=1}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}\idotsint\limits_{D_{n}}e^{i\mathbf{z}\cdot(\mathbf{x}_{1}+\cdots+\mathbf{x}_{n})}\frac{d\mathbf{x}_{1}\ldots d\mathbf{x}_{n}}{|\mathbf{x}_{1}|^{d}\cdot\ldots\cdot|\mathbf{x}_{n}|^{d}}\right),$
where $D_{n}\subset\mathbb{R}^{d\times n},$ $n\geq 1$ is defined as above. Notice first that the expression in the brackets above can be written as
$\int_{\mathbb{R}^{d}}e^{i\mathbf{z}\cdot\mathbf{x}}\left(\sum_{n=0}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}F^{(*n)}(\mathbf{x})\right),$
where (∗n) denotes the $n$-fold convolution and $F^{(*0)}(\cdot)=\delta_{\mathbf{0}}(\cdot)$. Moreover, the series above converges in $L_{2}$, indeed
$\biggl\|\sum_{n=0}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}F^{(*n)}(\mathbf{x})\biggl\|_{L_{2}}\leq$ $\sum_{n=0}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}\|F^{(*n)}\|_{L_{2}}$
$=$ $\sum_{n=0}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}\|\hat{F}\|_{L_{2}}^{n}=e^{-\theta\zeta_{d}\|\hat{F}\|_{L_{2}}}<\infty,$
where we have used the triangle inequality, the Plancherel theorem and the convolution property of the Fourier transform, see again [SW71].
Notice also that $|\mathbf{z}|^{-\theta}$ is the Fourier transform of a generalised (on a space of Schwartz functions) function $g(\mathbf{y})$ as defined above [GS69, p. 194].
Thus, it follows now from the convolution property of the Fourier transform (on a space of generalised functions this time) that $\hat{\mathscr{D}}(\mathbf{z})$ is the Fourier transform of a generalised function $f(\mathbf{x})$ as defined in (17). Here, we also used the fact that the space $L_{2}$ constitutes a subspace of the space of generalised functions.
Moreover, as noted in the discussion prior to the theorem, the distribution $\mathscr{D}(\frac{1}{\theta}I,\nu)$ is absolutely continuous. It means that there exists $q\in L_{1}(\mathbb{R}^{d})$, $q\geq 0$ such that
$\int\limits_{\mathbb{R}^{d}}q(\mathbf{x})d\mathbf{x}=1,$
$\int\limits_{\mathbb{R}^{d}}e^{i\mathbf{z}\cdot\mathbf{x}}q(\mathbf{x})d\mathbf{x}=\hat{\mathscr{D}}(\mathbf{z}).$
It follows then that the functions $f(\mathbf{x})$ and $q(\mathbf{x})$ should coincide in the weak sense, meaning
$\int_{\mathbb{R}^{d}}f(\mathbf{x})\phi(\mathbf{x})d\mathbf{x}=\int_{\mathbb{R}^{d}}q(\mathbf{x})\phi(\mathbf{x})d\mathbf{x}$
for every function $\phi(\mathbf{x})$ belonging to a space of test functions. Thus, it is not hard to show, using the relations above and choosing $0\leq\phi_{K}(\mathbf{x})\leq 1,$ $\mathbf{x}\in\mathbb{R}^{d}$ and $\phi_{K}(\mathbf{x})=1,$ $|\mathbf{x}|\leq K$, where $K>0$, that $\int_{\mathbb{R}^{d}}f(\mathbf{x})d\mathbf{x}=1$. The fact that $f(\mathbf{x})\geq 0$ for almost every (with respect to Lebesgue measure on $\mathbb{R}^{d}$) $\mathbf{x}\in\mathbb{R}^{d}$ follows similarly. We can then conclude that the equality $f(\mathbf{x})=q(\mathbf{x})$ holds in $L_{1}$ and, consequently, almost everywhere with respect to the Lebesgue measure on $\mathbb{R}^{d}$. Therefore, formula (17) gives an explicit expression for the density function of $\mathbf{X}$. ∎
Remark 7.
We note that formula (17) can be viewed as the Riesz potential of the $L_{2}$-function $\sum_{n=0}^{\infty}\frac{(-\theta\zeta_{d})^{n}}{n!}F^{(*n)}$, see e.g. [STE70, Chapter III].
Remark 8.
Note also that for $d=1$ formula (17) gives a formula for the convolution $f_{\theta/2}(x)*f_{\theta/2}(-x)$, $\theta>0,\theta\neq d+2n,n=0,1,\ldots$, where $f_{\theta/2}(x)$ is the density of the one-dimensional Dickman distribution $GD_{\theta/2}$. Moreover, it easily follows from [GR14, Formula 3.782(1)] that $\alpha_{1}=\gamma$, cf. Section 2.
Notice also that the restriction $\theta\neq d+2n,n=0,1,\ldots$ arises from the analytical techniques used in the proofs, rather than from any structural property of the distribution $\mathscr{D}(\frac{1}{\theta}I,\nu)$. In fact, the density $f(\mathbf{x})$ of $\mathscr{D}(\frac{1}{\theta}I,\nu)$ exists for any $\theta>0$, and the expression (17) remains valid except for values $\theta=d+2n,n\geq 0$.
The following proposition establishes the integro-partial differential equation for the density of distributions in $\mathscr{D}(\mathbb{I},H)$. It thus gives a multivariate counterpart for the equation (2), see also [CS75, Lemma 3]. Our result, however, holds in a weak sense, treating the density as a generalised function operating on the Schwartz space of rapidly decreasing functions [MIT13].
Proposition 3.16.
Let $\theta>0$ and $\nu\in H$ be given. Let $f(\mathbf{x})$, $\mathbf{x}\in\mathbb{R}^{d}$, denote the probability density function (given as a generalised function) of the $\mathscr{D}(\frac{1}{\theta}I,\nu)$ distribution. Then it satisfies the integro-partial differential equation
$\sum_{k=1}^{d}x_{k}\frac{\partial f(\mathbf{x})}{\partial x_{k}}=(\theta-d)f(\mathbf{x})-\theta\int\limits_{\mathbb{R}^{d}}f(\mathbf{x}-\mathbf{y})\nu(d\mathbf{y})$
in a weak sense.
Proof.
Notice that, due to Proposition 3.4, the Fourier transform $\hat{f}$ of $f$ satisfies
$\log\hat{f}(\mathbf{z})=\theta\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{is\mathbf{z}\cdot\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s},\;\mathbf{z}\in\mathbb{R}^{d}.$
Consequently,
$\sum_{k=1}^{\infty}z_{k}\frac{\partial\hat{f}(\mathbf{z})}{\partial z_{k}}=\theta\hat{f}(\mathbf{z})\hat{\nu}(\mathbf{z})-\theta\hat{f}(\mathbf{z}).$
The result follows now by properties of the Fourier transform on the space of generalised functions, see, for example, [MIT13, Th. 4.25 and Th. 4.33]. ∎
Remark 9.
The differential operator $L=\sum_{k=1}^{d}z_{k}\frac{\partial}{\partial z_{k}}$ is sometimes referred to as the Euler partial differential operator.
4 Some applications of the operator Dickman distributions
The one-dimensional Dickman distribution has applications in various probabilistic problems and beyond, as discussed in Section 2. Moreover, the multidimensional Dickman distribution from Definition 3.1 was proven to be useful in approximating small jumps of some multidimensional Lévy process. This section presents examples of how the above-mentioned applications can be extended to the case of operator Dickman distributions.
Hereafter $\overset{v}{\to}$” and $\overset{w}{\to}$” denote respectively vague and weak convergence of measures. In particular, $\mu_{n}\overset{v}{\to}\mu$, $n\to\infty$, for measures $\mu_{n}$ and $\mu$ over $\mathbb{R}^{d}$, whenever
$\int\limits_{\mathbb{R}^{d}}f(\mathbf{x})\mu_{n}(d\mathbf{x})\to\int\limits_{\mathbb{R}^{d}}f(\mathbf{x})\mu(d\mathbf{x}),\;n\to\infty,$
for all bounded, continuous maps $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ vanishing on a neighbourhood of $\mathbf{0}$ [KEN99, p. 41]. Moreover, by convention, $\overset{w}{\to}$”, when applied to random elements, means the weak convergence of their corresponding probability measures.
First, we present the following multidimensional counterpart for [VER72, Example 4.7.5].
Theorem 4.1.
Consider a sequence $\varepsilon_{1},\varepsilon_{2},\ldots$ of independent random variables such that $\mathbb{P}(\varepsilon_{k}=1)=1-\mathbb{P}(\varepsilon_{k}=0)=p_{k}$, $k\geq 1$, and assume that
$p_{k}\in[0,1),\;k\geq 1,$
$\lim\limits_{k\to\infty}p_{k}=0,\;\sum_{k=1}^{\infty}p_{k}=\infty.$
Let $L(n)$ denote the index of $n$-th 11 in the sequence $\varepsilon_{k},k\geq 1$.
Fix now $Q\in\mathscr{M}_{+}$ and assume that $Q$ is as in the Example 3.13. Then
$(L(n))^{Q}\sum_{k=1}^{\infty}(L(n+k))^{-Q}\mathbf{W}_{k}\overset{w}{\to}\mathbf{X},\text{ as }n\to\infty,$
where $\mathbf{W}_{1},\mathbf{W}_{2},\dots$ is the sequence of independent identically distributed random elements with a common law $\nu\in H$, independent also of a sequence $\varepsilon_{1},\varepsilon_{2},\ldots$, and $\mathscr{L}(\mathbf{X})=\mathscr{D}(Q,\nu)$.
Proof.
By the assumption $Q=S\Lambda S^{-1}$ for some invertible matrix $S$ and a diagonal matrix $\Lambda=diag\{a_{1},\ldots,a_{d}\}$, $a_{1},\ldots,a_{d}>0$. Then we have $t^{Q}=S\cdot diag\{t^{a_{1},},\ldots,t^{a_{d}}\}\cdot S^{-1}$ for any $t>0$. Since $a_{j}>0$, $j=1,\ldots,d$, it follows from [VER72, Ex. 4.7.5] that
$(L(n))^{a_{j}}\sum_{k=1}^{\infty}(L(n+k))^{-a_{j}}\overset{w}{\to}\sum_{k=1}^{\infty}(U_{1}\cdot\ldots\cdot U_{k})^{a_{j}}\text{ as }n\to\infty,$
for any $j=1,\ldots,d$. Rewrite now
$(L(n))^{Q}\sum_{k=1}^{\infty}(L(n+k))^{-Q}\mathbf{W}_{k}=S(L(n))^{\Lambda}\sum_{k=1}^{\infty}(L(n+k))^{-\Lambda}S^{-1}\mathbf{W}_{k}.$
The result follows then from the continuous mapping theorem [BIL99, Thm. 2.7], [BIL99, Thm. 2.8] and (8). ∎
Example 4.2.
Notice that the sequence $L(n),n\geq 1$, is an integer-valued Markov chain with transition probability function given by
$P(m,k)=\mathbb{P}(L(n)=k|L(n-1)=m)=p_{k}\prod_{j=m+1}^{k-1}(1-p_{j}),\;k\geq m+1,$
where $n\geq 1$ and $L(0)\coloneq 0$. Let us consider, for example, $p_{k}=\alpha/k,k\geq 1$ for some $\alpha\in(0,1)$. It then holds that
$\mathbb{P}(L(n)-L(n-1)$ $=r|L(n-1)=m)$
$=\left(1-\frac{\alpha}{m+1}\right)\cdot\ldots\cdot\left(1-\frac{\alpha}{m+r-1}\right)\frac{\alpha}{m+r},$
and we can recognise that, conditionally on L(n1)=mL(n-1)=m, the random variable L(n)L(n1)L(n)-L(n-1), $n\geq 1$, has generalised Sibuya distribution $S_{1}(\alpha,m)$ as defined in [KP18].
The sequence $L(n)$ has been studied extensively in relation to record epochs in [VER72].
Recall now that it has been proven in [COV09] that the Dickman distribution can be used to approximate small jumps of Lévy processes when the Brownian approximation is not applicable. It therefore makes simulation of the Lévy process (or the corresponding infinitely divisible random variable) easier. The analogous approximation for some class of multidimensional Lévy processes has been established in [GZ24a] using the multidimensional Dickman distribution defined as in (5). We show here that these results can be extended to a wider class of infinitely divisible distributions using the operator Dickman distributions $\mathscr{D}(Q,\sigma)$ with $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$. Note that our aim here is not to establish the most general case, but rather to demonstrate the possible scope of applications of the newly defined distributions.
We denote $\bar{\mathbb{R}}^{d}$ a one-point compactification of $\mathbb{R}^{d}$ and $\bar{\mathbb{R}}^{d}_{0}=\bar{\mathbb{R}}^{d}\setminus\{0\}$, $\bar{\mathbb{R}}_{+}=(0,\infty]$. We also introduce for any $Q\in\mathscr{M}_{+}$ the set $S_{Q}$ given by
$S_{Q}=\{\mathbf{x}\in\mathbb{S}^{d-1}:|r^{Q}\mathbf{x}|>1\text{ for every }r>1\},$
and a homeomorphic map
$\xi_{Q}:\bar{\mathbb{R}}_{+}\times S_{Q}\mapsto\bar{\mathbb{R}}^{d}_{0},\quad\xi_{Q}(r,\mathbf{x})=r^{Q}\mathbf{x},$
see the proof of [JUR82, Prop. 2].
To present the main result, we need the following generalisation of [GRA15, Lemma 4.9].
Lemma 4.3.
Let $\mu,\mu_{1},\mu_{2},\ldots$ be Borel measures on $\bar{\mathbb{R}}^{d}_{0}$. Assume also that $\mu$ is of the form
$\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{\bar{\mathbb{R}}_{+}}\mathds{1}_{B}(r^{Q}\mathbf{x})\sigma(d\mathbf{x})\lambda(dr),\;B\in\mathscr{B}(\bar{\mathbb{R}}^{d}_{0}),$ (19)
where $\sigma$ is a probability measure on $\mathbb{S}^{d-1}$ and $\lambda$ is a Borel measure on $\bar{\mathbb{R}}_{+}$ such that $\lambda\{|r|=a\}=0$ for any $0<a<\infty$ and $\lambda$ is finite on any set bounded away from 0. Then
$\mu_{n}\overset{v}{\to}\mu\quad\text{if and only if}\quad\mu_{n}(A_{t,D})\to\mu(A_{t,D}),\;n\to\infty$
for all $A_{t,D}=\xi_{Q}((t,\infty]\times D)=\{r^{Q}\mathbf{x}:r>t,\mathbf{x}\in D\}$ with $t>0$ and $D\in\mathscr{B}(S_{Q})$ such that $\sigma(\partial D)=0$.
Proof.
The proof repeats the proof of [GRA15, Lem. 4.9] almost word by word due to continuity and bijectivity of the map $\xi_{Q}$. ∎
Recall now that for $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$ the Lévy measure of the distribution $\mathscr{D}(Q,\sigma)$ is given by
$M(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}\mathds{1}_{B}(s^{Q}\mathbf{x})\sigma(d\mathbf{x})\frac{ds}{s},\;B\in\mathscr{B}(\mathbb{R}^{d}),$ (20)
which is of the form (19), see Corollary 3.6.
We will further limit our attention to $ID(\mathbb{R}^{d})$ distributions $\varrho$ with characteristic triple $(\mathbf{a},O,\mu)$ such that the Lévy measure $\mu$ satisfies the additional condition
$\int_{|\mathbf{x}|<1}|\mathbf{x}|\mu(d\mathbf{x})<\infty.$ (21)
Moreover, choosing for simplicity $\mathbf{a}=\int_{|\mathbf{x}|<1}\mathbf{x}\mu(d\mathbf{x})$, we obtain
$\hat{\varrho}(\mathbf{z})=\exp\left[\int\limits_{\mathbb{R}^{d}}(e^{i\mathbf{z}\cdot\mathbf{x}}-1)\mu(d\mathbf{x})\right].$ (22)
Theorem 4.4.
Fix $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$. Let $\varrho_{1},\varrho_{2},\ldots\in ID(\mathbb{R}^{d})$ be a sequence of infinitely divisible distributions satisfying the conditions (21) and (22). Let $\mu_{1},\mu_{2},\ldots$ denote the sequence of their Lévy measures.
Then $\varrho_{n}\overset{w}{\to}\mathscr{D}(Q,\sigma)$ as $n\to\infty$ whenever it holds for every $t>0$ and every $D\in\mathscr{B}(S_{Q})$, $\sigma(\partial D)=0$, that
$\mu_{n}(A_{t,D})\to\log(1\wedge t)\sigma(D),\;n\to\infty,$ (23)
where $A_{t,D}$ is given as in Lemma 4.3.
Proof.
Due to [KEN99, Th. 8.7] (see also [KAL21, Th. 7.7]) it is enough to check the following three conditions
  1. (i)
    $\mu_{n}\overset{v}{\to}M$;
  2. (ii)
    $\lim\limits_{\delta\downarrow 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})=0$ for any $\mathbf{z}\in\mathbb{R}^{d}$;
  3. (iii)
    $\lim\limits_{n\to\infty}\int\limits_{\mathbb{R}^{d}}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})=\int\limits_{\mathbb{R}^{d}}\mathbf{x}c(\mathbf{x})M(d\mathbf{x})$ for some bounded continuous function $c:\mathbb{R}^{d}\mapsto\mathbb{R}$ satisfying conditions of [KEN99, Th. 8.7].
It follows easily from Lemma 4.3 and (20) that conditions (i) and (23) are equivalent in the case of the $\mathscr{D}(Q,\sigma)$ distribution. Notice next that
$\lim\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq 1}|\mathbf{x}|\mu_{n}(d\mathbf{x})\leq\lim\limits_{n\to\infty}\int\limits_{\mathbb{R}^{d}}(|\mathbf{x}|\wedge 1)\mu_{n}(d\mathbf{x})=\int\limits_{\mathbb{R}^{d}}(|\mathbf{x}|\wedge 1)M(d\mathbf{x})<\infty,$
where we used the definition of vague convergence in the second equality and Corollary 3.6 on the last step. Thus, assuming that (i) holds, we can easily obtain (ii). Indeed, we have
$\lim\limits_{\delta\downarrow 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})\leq|\mathbf{z}|^{2}\lim\limits_{\delta\downarrow 0}\delta\cdot\int\limits_{|\mathbf{x}|\leq 1}|\mathbf{x}|M(d\mathbf{x})=0.$
Since for each $k=1,\ldots,d$ the component $x_{k}c(\mathbf{x})$ defines the bounded and continuous map $\mathbb{R}^{d}\mapsto\mathbb{R}$, which vanishes on a neighbourhood of $\mathbf{0}$ (see, in particular, conditions [KEN99, (8.3), (8.4)]), the condition (iii) follows directly from (i) and the definition of vague convergence. ∎
The next corollary provides a tractable example of the sequence $\mu_{n}$ (and therefore corresponding $ID$ distributions) satisfying the conditions of Theorem 4.4. It is a direct generalisation of [GZ24a, Cor. 1].
Corollary 4.5.
Let $\varrho_{\varepsilon}$, $\varepsilon>0$ be a collection of $ID(\mathbb{R}^{d})$ measures satisfying (22) and assume that the corresponding Lévy measures $\mu_{\varepsilon}$, $\varepsilon>0$, are of the form
$\mu_{\varepsilon}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varepsilon\rho(\varepsilon r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}).$
Assume also that for some $h:\mathbb{S}^{d-1}\mapsto\mathbb{R}$, $h\geq 0$, it holds that
$\int\limits_{\mathbb{S}^{d-1}}h(\mathbf{x})\sigma(d\mathbf{x})=b<\infty,\quad\lim\limits_{r\downarrow 0}\int\limits_{\mathbb{S}^{d-1}}|r\rho(r,\mathbf{x})-h(\mathbf{x})|\sigma(d\mathbf{x})=0.$
For simplicity, assume additionally that $b=1$. Then we have
$\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\tilde{\sigma})\;\text{ as }\varepsilon\downarrow 0,$
where $\tilde{\sigma}(B)=\int\limits_{B}h(\mathbf{x})\sigma(d\mathbf{x}),B\in\mathscr{B}(\mathbb{S}^{d-1})$.
Remark 10.
Let us consider the $ID(\mathbb{R}^{d})$ measure $\varrho$ satisfying (22) with Lévy measure $\tilde{\mu}$ given as
$\tilde{\mu}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}).$
For the measure $\tilde{\mu}$ and some $u>0$ we consider the truncation
$T_{u}\tilde{\mu}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{u}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
and note that $T_{u}\tilde{\mu}$ is a Lévy measure for any $u>0$. Note also that the truncation $T_{u}$ corresponds to considering only small jumps of the underlying Poisson random measure, whereas $\tilde{\mu}-T_{u}\tilde{\mu}$ corresponds to the compound Poisson part. Denote then $T_{u}\varrho$ the $ID(\mathbb{R}^{d})$ distribution with Lévy measure $T_{u}\tilde{\mu}$. Then $\varrho_{\varepsilon}$ of Corollary 4.5 can be obtained by the transformation $\varrho\mapsto\varepsilon^{-Q}T_{\varepsilon}\varrho$, which accords with truncation and operator rescaling of the original measure. Similar constructions have been considered in more detail in [COV09] and [GZ24a]. The application of this result for simulations has also been discussed there.
Note that we can also obtain the following simple corollary from Theorem 4.4, which is an extension of Proposition 3.8iii.
Corollary 4.6.
[…]
Proof.
The proof is analogous to the proof of 3.8iii if we notice that
$\frac{c_{1}+\cdots+c_{k}}{c}\frac{c_{1}\nu_{1}(\cdot)+\cdots+c_{k}\nu_{k}(\cdot)}{c_{1}+\cdots+c_{k}}\overset{v}{\to}\tilde{\nu}\;\text{as}\;k\to\infty$
by Lemma 4.3 and use Theorem 4.4. ∎
5 Simulations from the operator Dickman distribution
The random sample from the operator Dickman distribution can be simulated using the algorithm presented at the end of this section, where we use the representation (8). Moreover, the sum in (8) is truncated at $N$-th term, where $N=N_{max}\wedge K$ and $K=\inf\{k\in\mathbb{N}:\|(U_{1}\cdot\ldots\cdot U_{k})^{Q}\|_{\infty}<\epsilon\}$ for some fixed thresholds $\epsilon>0$ and $N_{max}\in\mathbb{N}$. Here $\|Q\|_{\infty}$ denotes the largest absolute value among the entries of $Q\in\mathscr{M}_{+}$.
In Figure 1, we show random samples of size 500 from the 2-dimensional operator Dickman distribution $\mathscr{D}(Q,\nu)$ with different operators $Q\in\mathscr{M}_{+}$ and $\nu$ being the uniform distribution on $\mathbb{S}^{1}$. The pictures demonstrate that the distribution $\mathscr{D}(Q,\nu)$ is rotationally invariant if $Q\in\mathbb{I}$ and that the covariance matrix of a random vector from the $\mathscr{D}(Q,\nu)$ distribution depends on the matrix $Q$, cf. Proposition 3.7 and Proposition 3.8 .
Figure 1: The random samples of size 500 from the 2-dimensional operator Dickman distribution $\mathscr{D}(Q,\nu)$ with $\nu$ being the uniform distribution on the $\mathbb{S}^{1}$ and various $Q$.
In Figure 2, we depict the first and second coordinates of 500 random $\mathbb{R}^{d}$-valued vectors sampled from $\mathscr{D}(I,\nu)$, $\nu$ being the uniform distribution on the unit sphere $\mathbb{S}^{d-1}$ for $d=3,6,9$. We can see that the coordinates of the random samples are concentrated closer to the origin for bigger values of $d$.
Figure 2: The first and second coordinates of random samples of size 500 from the operator Dickman distribution $\mathscr{D}(I,\nu)$ on $\mathbb{R}^{d}$, where $\nu$ is the uniform distribution on the unit sphere $\mathbb{S}^{d-1}$, for $d=3,6,9$.
Recall now that the von Mises distribution $\nu_{\kappa,\mu}$ on $\mathbb{S}^{1}$ is given as the distribution of the vector $(\cos\theta,\sin\theta)$ with the distribution of $\theta\in[0,2\pi)$ given via the probability density function $f_{\kappa,\mu}(x)$
$f_{\kappa,\mu}(x)=\frac{\exp(\kappa\cos(x-\mu))}{2\pi I_{0}(\kappa)},x\in[0,2\pi).$
Here $I_{0}(\cdot)$ is the modified Bessel function of the first kind of order 0, and $\kappa>0$, $\mu\in[0,2\pi)$ are the parameters. In Figure 3, we show the random samples of size 500 from the $2$-dimensional Dickman distribution $\mathscr{D}(I,\nu_{\kappa,\mu})$ with various values of parameters $\mu$ and $\kappa$. We can see that the points are distributed along the direction specified by the angle $\mu$, and the dispersion of points around this direction depends on $\kappa$.
Figure 3: The random samples of size 500 from the $2$-dimensional Dickman distribution $\mathscr{D}(I,\nu_{\mu,\kappa})$ with various values of parameters $\mu$ and $\kappa$.
Algorithm 1 Sampling from the operator Dickman distribution
1:$n$, $Q\in\mathscr{M}_{+}$, distribution $\nu$ on $\mathbb{R}^{d}$, tolerance $\varepsilon>0$, cap $N_{\max}$
[…]
3:function rDickman($n,Q,\epsilon,N_{\max}$)
4:  for $s=1$ to $n$ do
5:   $x\leftarrow 0_{d}$, $M\leftarrow I_{d}$
6:   repeat
7:     $u\sim U[0,1]$
8:     $M\leftarrow$ $e^{Q\log u}\cdot M$
9:     $x\leftarrow x+MW$, with $W\sim\nu$
10:     $terms\leftarrow terms+1$
11:   until $\|M\|_{\infty}<\epsilon$ or terms $N_{\max}$
12:   $X^{(s)}\leftarrow x$
13:  end for
14:  return $\{X^{(s)}\}$
15:end function
Acknowledgements
Nikolai Leonenko (NL) would like to thank for support the ARC Discovery Grant DP220101680 (Australia), Croatian Scientific Foundation (HRZZ) grant “Scaling in Stochastic Models” (IP-2022-10-8081), grant FAPESP 22/09201-8 (Brazil) and the Taith Research Mobility grant (Wales, Cardiff University).

### v2

@@
 2 One-dimensional Dickman distribution
-The one-dimensional Dickman distribution with parameter $\theta>0$ can be defined as a distribution of a random variable $X_{D}$ satisfying the equation
+The one-dimensional Dickman distribution with parameter $\theta>0$ can be defined as the distribution of a random variable $X_{D}$ satisfying the equation
 $X_{D}\overset{d}{=}U^{1/\theta}(1+X^{\prime}_{D}),$ (1)
@@
 Equivalently, the random variable $X_{D}$ is the fixed point (in a distributional sense) of the random affine transformation $\cdot\mapsto U^{1/\theta}(1+\cdot)$, where $U$ is as above.
-We will further denote $GD_{\theta}$ the Dickman distribution with parameter $\theta>0$.
+We will further denote $GD_{\theta}$, the Dickman distribution with parameter $\theta>0$.
 The distribution $GD_{\theta}$ is absolutely continuous for all $\theta>0$, and its density can be represented as
@@
 The density $f_{\theta}(x)$ of the $GD_{\theta}$ distribution has been an object of independent studies, and we recall here a few results related to it.
-The asymptotic behaviour of the density $f_{\theta}(x)$ for $\theta=1$ has been established in [DE 51b] and for general $\theta$ in [VER72, Lem.
+The asymptotic behaviour of the density $f_{\theta}(x)$ for $\theta=1$ has been established in [DE 51b] and for general $\theta>0$ in [VER72, Lem.
 4.7.9].
@@
 Let $\|Q\|\coloneq\sup_{|\mathbf{x}|\leq 1}|Q\mathbf{x}|$ denote the operator norm of $Q\in\mathscr{M}$.
-For any measure $\vartheta$ on $\mathbb{R}^{d}$ and any invertible operator $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ denote $T\vartheta$ the measure given by $T\vartheta(A)\coloneq\vartheta(T^{-1}(A))$, $A\in\mathscr{B}(\mathbb{R}^{d})$.
+For any measure $\vartheta$ on $\mathbb{R}^{d}$ and any measurable map $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, denote $T\vartheta$ the measure given by $T\vartheta(A)\coloneq\vartheta(T^{-1}(A))$, $A\in\mathscr{B}(\mathbb{R}^{d})$.
 Let $\mathscr{P}(\mathbb{R}^{d})$ denote the set of probability measures over $\mathbb{R}^{d}$ and for any $\pi\in\mathscr{P}(\mathbb{R}^{d})$ let $\hat{\pi}$ denote its characteristic function, i.e.
@@
 Note that $\mathscr{S}\subset H$.
-For a random vector $\mathbf{X}\in\mathbb{R}^{d}$ let $\mathscr{L}(\mathbf{X})\in\mathscr{P}(\mathbb{R}^{d})$ denote its probability distribution.
+For a random vector $\mathbf{X}\in\mathbb{R}^{d}$, let $\mathscr{L}(\mathbf{X})\in\mathscr{P}(\mathbb{R}^{d})$ denote its probability distribution.
 We also denote $ID(\mathbb{R}^{d})\subset\mathscr{P}(\mathbb{R}^{d})$ the set of all infinitely divisible distributions.
@@
 where $\theta>0$, $\mathbf{D}\overset{d}{=}\mathbf{D}^{\prime}$, $U$ has the uniform distribution on $[0,1]$, $\mathscr{L}(\mathbf{V})=\sigma\in\mathscr{S}$, and the random elements $U,\mathbf{V},\mathbf{D}^{\prime}$ are mutually independent.
-Next, we generalise this definition to a wider class of distributions by substituting the scalar function $u^{1/\theta},\theta>0$ with a matrix exponential $u^{Q},Q\in\mathscr{M}_{+}$ and a probability measure $\sigma\in\mathscr{S}$ with a probability measure $\nu\in H$.
+Next, we generalise this definition to a wider class of distributions by substituting the scalar function $u^{1/\theta},\theta>0$, with a matrix exponential $u^{Q},Q\in\mathscr{M}_{+}$, and a probability measure $\sigma\in\mathscr{S}$ with a probability measure $\nu\in H$.
 Definition 3.2.
@@
 where $\mathbf{X}\overset{d}{=}\mathbf{X}^{\prime}$, $U$ is uniformly distributed over $[0,1]$, $\mathscr{L}(\mathbf{W})=\nu$, and the random elements $U,\mathbf{X}^{\prime}$ and $\mathbf{W}$ are mutually independent.
-Moreover, for any $\mathscr{Q}\subseteq\mathscr{M}_{+}$ and $\mathscr{H}\subseteq H$ let $\mathscr{D}(\mathscr{Q},\mathscr{H})$ be a set of distributions defined as
+Moreover, for any $\mathscr{Q}\subseteq\mathscr{M}_{+}$ and $\mathscr{H}\subseteq H$, let $\mathscr{D}(\mathscr{Q},\mathscr{H})$ be a set of distributions defined as
 $\mathscr{D}(\mathscr{Q},\mathscr{H})=\{\mathscr{D}(Q,\nu),Q\in\mathscr{Q},\nu\in\mathscr{H}\},$
@@
 6] showed that the limit exists if $\infty<\mathbb{E}\log\|M_{1}\|<0$ and $0<\mathbb{E}|\bm{\xi}_{1}|^{\kappa}<\infty$ for some $\kappa>0$.
-However, as shown in [VER72] for $d=1$, the last condition can be weakened to $\mathbb{E}\log^{+}|\xi_{1}|<\infty$ under the assumption that the first condition holds, and this proof translates easily for $d>1$.
+However, as shown in [VER79] for $d=1$, the last condition can be weakened to $\mathbb{E}\log^{+}|\xi_{1}|<\infty$ under some assumptions on $M_{1}$, and this proof translates easily for $d>1$ in our case, using (11) below.
 Hereafter, we define
@@
 We shall exclude it from consideration and, furthermore, assume that $\nu(\{\mathbf{0}\})=0$.
+Note also that throughout the work we will make use of the following estimate.
+Let $Q\in\mathscr{M}_{+}$ be fixed, then there exist positive constants $c_{1},c_{2}$ and $k,K$ such that for all $0\leq s\leq 1$ it holds that
+$c_{1}s^{K}\leq\|s^{Q}\|\leq c_{2}s^{k},$ (11)
+see [MS01, Prop. 2.2.11] and the proof of [URB72, Lem. 6.1].
 To establish properties of the operator Dickman distributions, we will need the following lemma.
@@
 Proof.
-(i) First note that the function $f(t)=t\wedge 1$ is non-decreasing.
-Note also that for any $Q\in\mathscr{M}_{+}$ there exist $K>0$ and $c_{1}>0$ such that $|s^{Q}\mathbf{x}|\leq c_{1}s^{K}|\mathbf{x}|$ for all $s>0$ and $\mathbf{x}\in\mathbb{R}^{d}$ [MS01, Prop.
-2.2.11], see also the proof of [URB72, Lem.
-6.1.].
-We might as well choose $c_{1}>1$, thus
-$\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|\wedge 1)\frac{ds}{s}$ $\leq c_{1}\int\limits_{0}^{1}(s^{K}|\mathbf{x}|\wedge 1)\frac{ds}{s}=\frac{c_{1}}{K}\int\limits_{0}^{1}(s|\mathbf{x}|\wedge 1)\frac{ds}{s}$
-$=\frac{c_{1}}{K}\int\limits_{0}^{|\mathbf{x}|}(s\wedge 1)\frac{ds}{s}=\frac{c_{1}}{K}\left(1\wedge|\mathbf{x}|-\log\left(1\wedge|\mathbf{x}|^{-1}\right)\right),$
+(i) First note that the function $f(t)=t\wedge 1$ is non-decreasing. We can assume, without loss of generality, that $c_{2}>1$ in (11). Thus, we obtain
+$\int\limits_{0}^{1}(|s^{Q}\mathbf{x}|\wedge 1)\frac{ds}{s}$ $\leq c_{2}\int\limits_{0}^{1}(s^{k}|\mathbf{x}|\wedge 1)\frac{ds}{s}=\frac{c_{2}}{k}\int\limits_{0}^{1}(s|\mathbf{x}|\wedge 1)\frac{ds}{s}$
+$=\frac{c_{2}}{k}\int\limits_{0}^{|\mathbf{x}|}(s\wedge 1)\frac{ds}{s}=\frac{c_{2}}{k}\left(1\wedge|\mathbf{x}|-\log\left(1\wedge|\mathbf{x}|^{-1}\right)\right),$
 and the result follows.
@@
 $\log\psi(r^{Q^{*}}\bm{\omega})=\int\limits_{0}^{r}(\hat{\nu}(s^{Q^{*}}\bm{\omega})-1)\frac{ds}{s}+g(\bm{\omega})=\int\limits_{0}^{1}(\hat{\nu}((sr)^{Q^{*}}\bm{\omega})-1)\frac{ds}{s}+g(\bm{\omega}).$
-Letting $r=0$ we can deduce from $\psi(\mathbf{0})=1$ that $g(\bm{\omega})=0$ for all $\bm{\omega}\in\mathbb{R}^{d}$.
+Letting $r=0$, we can deduce from $\psi(\mathbf{0})=1$ that $g(\bm{\omega})=0$ for all $\bm{\omega}\in\mathbb{R}^{d}$.
 Returning to the variable $\mathbf{z}$, we finally get
@@
 Remark 4.
-Remarkably, the distributions with Lévy measure as in (12) have previously appeared in [JUR82].
+Remarkably, the distributions with Lévy measure as in (13) previously appeared in [JUR82].
 Proposition 3.7.
   1. (i)
-    Assume that $\nu\in H$ is such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(\mathbf{x})<\infty$.
-Then the mean vector $\mathbf{m}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and is given as
+    Assume that $\nu\in H$ is such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty$.
+Then the mean vector $\mathbf{m}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and can be given by
     $\mathbf{m}_{Q,\nu}=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}s^{Q}\mathbf{x}\nu(d\mathbf{x})\frac{ds}{s}.$ (14)
   2. (ii)
-    Assume that $\nu\in H$ such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(\mathbf{x})<\infty$.
+    Assume that $\nu\in H$ such that $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(d\mathbf{x})<\infty$.
 Then the covariance matrix $\mathbb{C}_{Q,\nu}$ of $\mathscr{D}(Q,\nu)$ exists for any $Q\in\mathscr{M}_{+}$ and can be given by
@@
 Both formulas (14) and (15) follow easily by differentiation of the characteristic function (12).
-Note that the finiteness of the integrals in (13) under the assumptions of (i) follows again by the estimate
-$|s^{Q}\mathbf{x}|\leq{c_{1}}s^{K}|\mathbf{x}|,\;s>0,\mathbf{x}\in\mathbb{R}^{d}$
-with some $c_{1},K>0$. Consequently, we obtain
-$|\mathbf{m}|\leq\frac{c_{1}}{K}\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty.$
+Note that the finiteness of the integrals in (14) under the assumptions of (i) follows again from (11).
+Indeed, we can obtain
+$|\mathbf{m}|\leq\frac{c_{2}}{k}\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|\nu(d\mathbf{x})<\infty.$
 The finiteness of the integrals in (15) follows similarly. ∎
@@
 and integration by parts, that formulas (14) and (15) can be rewritten as
-$\begin{gathered}\mathbf{m}_{Q,\nu}=Q^{-1}\int\limits_{\mathbb{R}^{d}}\mathbf{x}\nu(d\mathbf{x})\;\text{ and }\;\mathbb{C}_{Q,\nu}=\mathscr{A}_{Q}^{-1}\left(\int\limits_{\mathbb{R}^{d}}\mathbf{x}\mathbf{x}^{T}\nu(d\mathbf{x})\right),\end{gathered}$
+$\begin{gathered}\mathbf{m}_{Q,\nu}=Q^{-1}\int_{\mathbb{R}^{d}}\mathbf{x}\nu(d\mathbf{x})\;\text{ and }\;\mathbb{C}_{Q,\nu}=\mathscr{A}_{Q}^{-1}\left(\int_{\mathbb{R}^{d}}\mathbf{x}\mathbf{x}^{T}\nu(d\mathbf{x})\right),\end{gathered}$
 where $\mathscr{A}_{Q}:\mathscr{M}_{d}\mapsto\mathscr{M}_{d}$, $\mathscr{A}_{Q}(B)=QB+BQ^{*}$.
@@
   3. (iii)
-    For any fixed $Q\in\mathscr{M}_{+}$ the class $\mathscr{D}(\frac{1}{c}Q,H)$, where $c$ runs over $(0,\infty)$, is closed under finite convolutions.
+    For any fixed $Q\in\mathscr{M}_{+}$, the class $\mathscr{D}(w^{-1}Q,H)$, where $w$ runs over $(0,\infty)$, is closed under finite convolutions.
 Proof.
@@
 The proof of part (ii) is analogous.
-(iii) Note first that for fixed $Q\in\mathscr{M}_{+}$, $c>0$ and $\mathscr{D}=\mathscr{D}(\frac{1}{c}Q,\nu)$, from (11) by a change of variables we obtain
-$\log\widehat{\mathscr{D}}(\mathbf{z})=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{\frac{1}{c}Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}=c\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}.$
-Let $Q\in\mathscr{M}_{+}$, $c_{1},c_{2}>0$ and $\nu_{1},\nu_{2}\in H$ and consider the independent distributions $\mathscr{D}_{1}=\mathscr{D}(\frac{1}{c_{1}}Q,\nu_{1})$, $\mathscr{D}_{2}=\mathscr{D}(\frac{1}{c_{2}}Q,\nu_{2})$.
+(iii) Note first that for fixed $Q\in\mathscr{M}_{+}$, $w>0$ and $\mathscr{D}=\mathscr{D}(w^{-1}Q,\nu)$, from (12) by a change of variables we obtain
+$\log\widehat{\mathscr{D}}(\mathbf{z})=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{\frac{1}{w}Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}=w\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\nu(d\mathbf{x})\frac{ds}{s}.$
+Let $Q\in\mathscr{M}_{+}$, $w_{1},w_{2}>0$ and $\nu_{1},\nu_{2}\in H$ and consider the independent distributions $\mathscr{D}_{1}=\mathscr{D}(w_{1}^{-1}Q,\nu_{1})$, $\mathscr{D}_{2}=\mathscr{D}(w_{2}^{-1}Q,\nu_{2})$.
 If we denote $\mathscr{D}=\mathscr{D}_{1}*\mathscr{D}_{2}$, then
-$\log\widehat{\mathscr{D}}(\mathbf{z})$ $=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)(c_{1}\nu_{1}+c_{2}\nu_{2})(d\mathbf{x})\frac{ds}{s}$
-$=(c_{1}+c_{2})\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\frac{(c_{1}\nu_{1}+c_{2}\nu_{2})(d\mathbf{x})}{c_{1}+c_{2}}\frac{ds}{s},$
-and it is easy to see that $\mathscr{D}=\mathscr{D}(\frac{1}{c_{1}+c_{2}}Q,\frac{c_{1}\nu_{1}+c_{2}\nu_{2}}{c_{1}+c_{2}})$. ∎
+$\log\widehat{\mathscr{D}}(\mathbf{z})$ $=\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)(w_{1}\nu_{1}+w_{2}\nu_{2})(d\mathbf{x})\frac{ds}{s}$
+$=(w_{1}+w_{2})\int\limits_{\mathbb{R}^{d}}\int\limits_{0}^{1}(e^{i\mathbf{z}\cdot s^{Q}\mathbf{x}}-1)\frac{(w_{1}\nu_{1}+w_{2}\nu_{2})(d\mathbf{x})}{w_{1}+w_{2}}\frac{ds}{s},$
+and it is easy to see that $\mathscr{D}=\mathscr{D}(\frac{1}{w_{1}+w_{2}}Q,\frac{w_{1}\nu_{1}+w_{2}\nu_{2}}{w_{1}+w_{2}})$. ∎
 Remark 6.
@@
 Let $\mathscr{L}(\mathbf{X})=\mathscr{D}(\Lambda,\nu)$ for some $\nu\in H$.
-Since in this case it holds for $t>0$ that $t^{\Lambda}=diag(s^{a_{1}},\ldots,s^{a_{d}})$ it is clear that the components $X_{i}$ of the vector $\mathbf{X}$ satisfy the distributional equation
+Since in this case it holds for $s>0$ that $s^{\Lambda}=diag(s^{a_{1}},\ldots,s^{a_{d}})$, it is clear that the components $X_{i}$ of the vector $\mathbf{X}$ satisfy the distributional equation
 $X_{i}\overset{d}{=}U^{a_{i}}(X_{i}+W_{i}),\;i=1,\ldots,d,$
@@
 Moreover, the formula (15) for the covariance matrix in this case simplifies to
-$C_{Q,\nu}]_{i,j}=\frac{1}{a_{i}+a_{j}}\int\limits_{\mathbb{R}^{d}}x_{i}x_{j}\nu(d\mathbf{x}),\;i,j=1\ldots,d$
-whenever $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(\mathbf{x})<\infty$.
+$\mathbf{C}_{Q,\nu}]_{i,j}=\frac{1}{a_{i}+a_{j}}\int\limits_{\mathbb{R}^{d}}x_{i}x_{j}\nu(d\mathbf{x}),\;i,j=1\ldots,d$
+whenever $\int\limits_{\mathbb{R}^{d}}|\mathbf{x}|^{2}\nu(d\mathbf{x})<\infty$.
 Notice also that in the simplest case $\nu=\delta_{\mathbf{w}}$, for some $\mathbf{w}=(w_{1},\ldots,w_{d})\in\mathbb{R}^{d}_{+}$, the components $w_{i}X_{i}$ are distributed according to $GD_{1/a_{i}}$, $i=1,\ldots,d$.
@@
 Example 3.14.
-Consider $\mathscr{D}=\mathscr{D}(\frac{1}{\theta}I,\nu)$, $\theta>0$ and assume $\nu\in\mathscr{S}$.
+Consider $\mathscr{D}=\mathscr{D}(\frac{1}{\theta}I,\nu)$, $\theta>0$, and assume $\nu\in\mathscr{S}$.
 As was noted before, the random vector $\mathbf{X}$ with $\mathscr{L}(\mathbf{X})=\mathscr{D}$ has the multidimensional Dickman distribution as defined in [BM20].
@@
 In this work, the explicit formula for the density is obtained for a specific case of the multidimensional Dickman distribution $\mathscr{D}\in\mathscr{D}(\mathbb{I},\mathscr{S})$ in Theorem 3.15, whereas Proposition 3.16 treats a more general case of $\mathscr{D}(\mathbb{I},H)$.
-Note that we utilise some elements of the theory of generalised functions operating on the Schwartz space of rapidly decreasing functions; for more details on this topic, see the classic book [GS69] and also [MIT13].
+Note that we utilise some elements of the theory of generalised functions operating on the Schwartz space of rapidly decreasing functions; for more details on this topic, see the classical book [GS69] and also [MIT13].
 We consider first the multidimensional Dickman distribution $\mathscr{D}(\frac{1}{\theta}I,\nu)$, where the distribution $\nu$ is uniform on the sphere $\mathbb{S}^{d-1}$.
@@
 4.7.7].
-We further denote $J_{a}(\cdot),a>-\frac{1}{2}$ the Bessel function of first kind, $Leb$ the Lebesgue measure on $\mathbb{S}^{d-1}$, $\zeta_{d}=(Leb(\mathbb{S}^{d-1}))^{-1}={\Gamma(d/2)}\cdot(2\pi)^{-d/2}$ the reciprocal surface area of $\mathbb{S}^{d-1}$ and $Y_{d}(s)=(s/2)^{1-\frac{d}{2}}\Gamma(d/2)J_{d/2-1}(s)$, $s>0$, $d>1$, the spherical Bessel function, see e.g.
+We further denote $J_{a}(\cdot),a>-\frac{1}{2}$, the Bessel function of first kind, $Leb$ the Lebesgue measure on $\mathbb{S}^{d-1}$, $\zeta_{d}=(Leb(\mathbb{S}^{d-1}))^{-1}={\Gamma(d/2)}\cdot(2\pi)^{-d/2}$ the reciprocal surface area of $\mathbb{S}^{d-1}$ and $Y_{d}(s)=(s/2)^{1-\frac{d}{2}}\Gamma(d/2)J_{d/2-1}(s)$, $s>0$, $d>1$, the spherical Bessel function, see e.g.
 [PIN93].
@@
 $=$ $e^{\theta\alpha_{d}}|\mathbf{z}|^{-\theta}\exp\left(-\theta\int\limits_{1}^{\infty}Y_{d}(s|\mathbf{z}|)\frac{ds}{s}\right).$
-Denote now $\tilde{F}_{k}(s)=s^{-d}\mathds{1}\{1<s<k\}$, $k\geq 2$ and $\tilde{F}(s)=s^{-d}\mathds{1}\{s>1\}$.
+Denote now $\tilde{F}_{k}(s)=s^{-d}\mathds{1}\{1<s<k\}$, $k\geq 2$, and $\tilde{F}(s)=s^{-d}\mathds{1}\{s>1\}$.
 Let also $F_{k}(\mathbf{x})=\tilde{F}_{k}(|\mathbf{x}|)$ and $F(\mathbf{x})=\tilde{F}(|\mathbf{x}|)$.
@@
 Consequently,
-$\sum_{k=1}^{\infty}z_{k}\frac{\partial\hat{f}(\mathbf{z})}{\partial z_{k}}=\theta\hat{f}(\mathbf{z})\hat{\nu}(\mathbf{z})-\theta\hat{f}(\mathbf{z}).$
+$\sum_{k=1}^{d}z_{k}\frac{\partial\hat{f}(\mathbf{z})}{\partial z_{k}}=\theta\hat{f}(\mathbf{z})\hat{\nu}(\mathbf{z})-\theta\hat{f}(\mathbf{z}).$
 The result follows now by properties of the Fourier transform on the space of generalised functions, see, for example, [MIT13, Th. 4.25 and Th. 4.33]. ∎
@@
 $P(m,k)=\mathbb{P}(L(n)=k|L(n-1)=m)=p_{k}\prod_{j=m+1}^{k-1}(1-p_{j}),\;k\geq m+1,$
-where $n\geq 1$ and $L(0)\coloneq 0$. Let us consider, for example, $p_{k}=\alpha/k,k\geq 1$ for some $\alpha\in(0,1)$. It then holds that
+where $n\geq 1$ and $L(0)\coloneq 0$. Let us consider, for example, $p_{k}=\alpha/k$, $k\geq 1$, for some $\alpha\in(0,1)$. It then holds that
 $\mathbb{P}(L(n)-L(n-1)$ $=r|L(n-1)=m)$
-$=\left(1-\frac{\alpha}{m+1}\right)\cdot\ldots\cdot\left(1-\frac{\alpha}{m+r-1}\right)\frac{\alpha}{m+r},$
+$=\left(1-\frac{\alpha}{m+1}\right)\cdot\ldots\cdot\left(1-\frac{\alpha}{m+r-1}\right)\frac{\alpha}{m+r},\;r\geq 1$
 and we can recognise that, conditionally on L(n1)=mL(n-1)=m, the random variable L(n)L(n1)L(n)-L(n-1), $n\geq 1$, has generalised Sibuya distribution $S_{1}(\alpha,m)$ as defined in [KP18].
@@
 Note that our aim here is not to establish the most general case, but rather to demonstrate the possible scope of applications of the newly defined distributions.
-We denote $\bar{\mathbb{R}}^{d}$ a one-point compactification of $\mathbb{R}^{d}$ and $\bar{\mathbb{R}}^{d}_{0}=\bar{\mathbb{R}}^{d}\setminus\{0\}$, $\bar{\mathbb{R}}_{+}=(0,\infty]$.
+We denote $\bar{\mathbb{R}}^{d}$, a one-point compactification of $\mathbb{R}^{d}$ and $\bar{\mathbb{R}}^{d}_{0}=\bar{\mathbb{R}}^{d}\setminus\{0\}$, $\bar{\mathbb{R}}_{+}=(0,\infty]$.
 We also introduce for any $Q\in\mathscr{M}_{+}$ the set $S_{Q}$ given by
@@
 The proof repeats the proof of [GRA15, Lem. 4.9] almost word by word due to continuity and bijectivity of the map $\xi_{Q}$. ∎
-Recall now that for $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$ the Lévy measure of the distribution $\mathscr{D}(Q,\sigma)$ is given by
+Recall now that for $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$, the Lévy measure of the distribution $\mathscr{D}(Q,\sigma)$ is given by
 $M(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}\mathds{1}_{B}(s^{Q}\mathbf{x})\sigma(d\mathbf{x})\frac{ds}{s},\;B\in\mathscr{B}(\mathbb{R}^{d}),$ (21)
@@
 Moreover, choosing for simplicity $\mathbf{a}=\int_{|\mathbf{x}|<1}\mathbf{x}\mu(d\mathbf{x})$, we obtain
-$\hat{\varrho}(\mathbf{z})=\exp\left[\int\limits_{\mathbb{R}^{d}}(e^{i\mathbf{z}\cdot\mathbf{x}}-1)\mu(d\mathbf{x})\right].$ (22)
+$\hat{\varrho}(\mathbf{z})=\exp\left[\int_{\mathbb{R}^{d}}(e^{i\mathbf{z}\cdot\mathbf{x}}-1)\mu(d\mathbf{x})\right].$ (23)
 Theorem 4.4.
 Fix $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$.
-Let $\varrho_{1},\varrho_{2},\ldots\in ID(\mathbb{R}^{d})$ be a sequence of infinitely divisible distributions satisfying the conditions (21) and (22).
-Let $\mu_{1},\mu_{2},\ldots$ denote the sequence of their Lévy measures.
+Let $\varrho_{1},\varrho_{2},\ldots\in ID(\mathbb{R}^{d})$ be a sequence of infinitely divisible distributions and let $\mu_{1},\mu_{2},\ldots$ denote the sequence of their Lévy measures, which we assume satisfy the condition (22).
+Assume additionally that
+$\lim_{\delta\to 0}\limsup_{n\to\infty}\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{n}(d\mathbf{x})=0.$ (24)
 Then $\varrho_{n}\overset{w}{\to}\mathscr{D}(Q,\sigma)$ as $n\to\infty$ whenever it holds for every $t>0$ and every $D\in\mathscr{B}(S_{Q})$, $\sigma(\partial D)=0$, that
-$\mu_{n}(A_{t,D})\to\log(1\wedge t)\sigma(D),\;n\to\infty,$ (23)
+$\mu_{n}(A_{t,D})\to-\log(1\wedge t)\sigma(D),\;n\to\infty,$ (25)
 where $A_{t,D}$ is given as in Lemma 4.3.
 Proof.
-Due to [KEN99, Th. 8.7] (see also [KAL21, Th. 7.7]) it is enough to check the following three conditions
+Due to [KEN99, Th. 8.7], it is enough to check the following three conditions
   1. (i)
@@
   2. (ii)
-    $\lim\limits_{\delta\downarrow 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})=0$ for any $\mathbf{z}\in\mathbb{R}^{d}$;
+    $\lim\limits_{\delta\to 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})=0$ for any $\mathbf{z}\in\mathbb{R}^{d}$;
   3. (iii)
@@
 8.7].
-It follows easily from Lemma 4.3 and (20) that conditions (i) and (23) are equivalent in the case of the $\mathscr{D}(Q,\sigma)$ distribution. Notice next that
-$\lim\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq 1}|\mathbf{x}|\mu_{n}(d\mathbf{x})\leq\lim\limits_{n\to\infty}\int\limits_{\mathbb{R}^{d}}(|\mathbf{x}|\wedge 1)\mu_{n}(d\mathbf{x})=\int\limits_{\mathbb{R}^{d}}(|\mathbf{x}|\wedge 1)M(d\mathbf{x})<\infty,$
-where we used the definition of vague convergence in the second equality and Corollary 3.6 on the last step. Thus, assuming that (i) holds, we can easily obtain (ii). Indeed, we have
-$\lim\limits_{\delta\downarrow 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})\leq|\mathbf{z}|^{2}\lim\limits_{\delta\downarrow 0}\delta\cdot\int\limits_{|\mathbf{x}|\leq 1}|\mathbf{x}|M(d\mathbf{x})=0.$
-Since for each $k=1,\ldots,d$ the component $x_{k}c(\mathbf{x})$ defines the bounded and continuous map $\mathbb{R}^{d}\mapsto\mathbb{R}$, which vanishes on a neighbourhood of $\mathbf{0}$ (see, in particular, conditions [KEN99, (8.3), (8.4)]), the condition (iii) follows directly from (i) and the definition of vague convergence.
+It follows from Lemma 4.3 and (21) that conditions (i) and (25) are equivalent in the case of the $\mathscr{D}(Q,\sigma)$ distribution.
+Notice also that the condition (ii) follows readily from (24).
+Consider condition (iii) now.
+Assuming that (i) holds, we can rewrite for any $\delta>0$
+$\lim\limits_{n\to\infty}\int\limits_{\mathbb{R}^{d}}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})=$ $\lim\limits_{n\to\infty}\left(\int\limits_{|\mathbf{x}|<\delta}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})+\int\limits_{|\mathbf{x}|\geq\delta}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})\right)$
+$=$ $\lim\limits_{n\to\infty}\int\limits_{|\mathbf{x}|<\delta}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})+\int\limits_{|\mathbf{x}|\geq\delta}\mathbf{x}c(\mathbf{x})M(d\mathbf{x}),$
+where the last equality follows from the definition of vague convergence since the maps $\mathbb{R}^{d}\mapsto\mathbb{R}$, $\mathbf{x}\mapsto x_{k}c(\mathbf{x})\mathds{1}{|\mathbf{x}|\geq\delta}$, $k=1,\ldots,d$, define bounded continuous functions, which vanish at the neighbourhood of $\mathbf{0}$.
+We can now let $\delta\to 0$ and (iii) follows from (24).
 ∎
-The next corollary provides a tractable example of the sequence $\mu_{n}$ (and therefore corresponding $ID$ distributions) satisfying the conditions of Theorem 4.4.
+We now consider the specific example of a sequence satisfying the conditions of Theorem 4.4. Namely let $\mu$ be a Lévy measure on $\mathbb{R}^{d}$ such that
+$\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varpi(d\mathbf{x},dr),\quad B\in\mathscr{B}(\mathbb{R}^{d}),$ (26)
+where $Q\in\mathscr{M}_{+}$ is fixed and $\varpi(\cdot,\cdot)$ is a measure on $\mathbb{S}^{d-1}\times\mathbb{R}_{+}$ such that
+$\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}|r^{Q}\mathbf{x}|\mathds{1}_{B_{1}}(r^{Q}\mathbf{x})\varpi(d\mathbf{x},dr)<\infty,$ (27)
+where we denote $B_{h}\subset\mathbb{R}^{d}$, an open ball of radius $h>0$. Thus, measure $\mu$ satisfies (22) and there exists $\rho\in ID(\mathbb{R}^{d})$ such that (23) holds.
+For the measure $\mu$ and some $u>0$, we consider the truncation
+$T_{u}\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{u}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varpi(d\mathbf{x},dr),\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
+and note that $T_{u}\mu$ is a Lévy measure for any $u>0$.
+Note also that the truncation $T_{u}$ corresponds to considering only small jumps of the underlying Poisson random measure, whereas the measure $\mu-T_{u}\mu$ corresponds to the compound Poisson part.
+Denote then $T_{u}\varrho$, the $ID(\mathbb{R}^{d})$ distribution with Lévy measure $T_{u}\mu$.
+Consider further the transformation $\varrho\mapsto\varepsilon^{-Q}T_{\varepsilon}\varrho$, $\varepsilon>0$, which accords with truncation and operator rescaling of the original measure and denote $\varrho_{\varepsilon}=\varepsilon^{-Q}T_{\varepsilon}\varrho$.
+Let $\mu_{\varepsilon}$ be the corresponding Lévy measure, i.e.
+$\mu_{\varepsilon}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\varepsilon}\mathds{1}_{B}\{(r/\varepsilon)^{Q}\mathbf{x}\}\varpi(d\mathbf{x},dr),\quad B\in\mathscr{B}(\mathbb{R}^{d}).$ (28)
+We show further that for measures $\mu_{\varepsilon}$, $\varepsilon\to 0$, the condition (24) follows from (25).
+Consequently, the vague convergence of the Lévy measures $\mu_{\varepsilon}\overset{v}{\to}M$ is sufficient to establish weak convergence $\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\sigma)$, $\varepsilon\to 0$, for a sequence $\varrho_{\varepsilon}$ obtained by the transformation of measure $\varrho$.
+We note that similar constructions have been considered in more detail in [COV09] and [GZ24a].
+The application of this result for simulations has also been discussed there.
+Lemma 4.5.
+Let $\mu_{\varepsilon}$, $\varepsilon>0$, be a collection of Lévy measures given by (28).
+Assume also that there exists some probability measure $\sigma$ on $\mathbb{S}^{d-1}$ such that for every $t>0$ and every $D\in\mathscr{B}(S_{Q})$, $\sigma(\partial D)=0$, it holds that
+$\mu_{\varepsilon}(A_{t,D})\to-\log(1\wedge t)\sigma(D),\;\varepsilon\to 0,$
+where $A_{t,D}$ is given as in Lemma 4.3. Then $\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\sigma)$.
+Proof.
+Note the assumption above implies that $\mu_{\varepsilon}\overset{v}{\to}M$, where measure $M$ is given as in (21), see Lemma 4.3. We will now show that
+$\lim_{\delta\to 0}\limsup_{\varepsilon\to 0}\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})=0.$ (29)
+The idea of the proof is due to [GZ24a, Prop. 1]. Rewrite now for some $\delta,\varepsilon>0$
+$\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})=$ $\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\varepsilon}|(r/\varepsilon)^{Q}\mathbf{x}|\mathds{1}_{B_{\delta}}((r/\varepsilon)^{Q}\mathbf{x})\varpi(d\mathbf{x},dr)$
+$=$ $\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}|r^{Q}\mathbf{x}|\mathds{1}_{B_{\delta}}(r^{Q}\mathbf{x})\varpi(d\mathbf{x},\varepsilon dr).$
+Using (11), we get $|r^{Q}\mathbf{x}|\geq c_{1}r^{K}|\mathbf{x}|=c_{1}r^{K}$ for all $0\leq r\leq 1$ and some $c_{1},K>0$. Denote then $g(\delta)=(\delta/c_{1})^{1/K}$ and notice that
+$\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})\leq$ $\int_{\mathbb{S}^{d-1}}\int_{0}^{g(\delta)}|r^{Q}\mathbf{x}|\varpi(d\mathbf{x},\varepsilon dr)$
+$=$ $\sum_{n=1}^{\infty}\int_{\mathbb{S}^{d-1}}\int_{g(\delta)2^{-n}}^{g(\delta)2^{-n+1}}|r^{Q}\mathbf{x}|\varpi(d\mathbf{x},\varepsilon dr)$
+$=$ $\sum_{n=1}^{\infty}\int_{\mathbb{S}^{d-1}}\int_{1/2}^{1}\left|\left(\frac{rg(\delta)}{2^{n-1}}\right)^{Q}\mathbf{x}\right|\varpi\left(d\mathbf{x},\varepsilon\frac{g(\delta)}{2^{n-1}}dr\right)$
+$\leq$ $\sum_{n=1}^{\infty}c_{2}\left(\frac{g(\delta)}{2^{n-1}}\right)^{k}\int_{\mathbb{S}^{d-1}}\int_{1/2}^{1}\left|r^{Q}\mathbf{x}\right|\varpi\left(d\mathbf{x},\varepsilon\frac{g(\delta)}{2^{n-1}}dr\right),$
+where we applied (11) with $s=g(\delta)/2^{n-1}<1$, assuming that $\delta$ is small enough. Notice now that
+$\int_{\mathbb{S}^{d-1}}\int_{1/2}^{1}\left|r^{Q}\mathbf{x}\right|\varpi\left(d\mathbf{x},\varepsilon\frac{g(\delta)}{2^{n-1}}dr\right)=\int_{\mathbb{R}^{d}}|\mathbf{y}|\mathds{1}_{A_{1/2}^{1}}(\mathbf{y})\mu_{\varepsilon\frac{g(\delta)}{2^{n-1}}}(d\mathbf{y}),$
+where $A_{a}^{b}=\{\mathbf{y}\in\mathbb{R}^{d}\,|\mathbf{y}=r^{Q}\mathbf{x},r\in(a,b),\mathbf{x}\in\mathbb{S}^{d-1}\}$, $0\leq a<b<\infty$.
+By the definition of vague convergence, the last integral converges as $\varepsilon\to 0$ to
+$\int_{\mathbb{R}^{d}}|\mathbf{y}|\mathds{1}_{A_{1/2}^{1}}(\mathbf{y})M(d\mathbf{y})=m^{*}<\infty.$
+Consequently, for any $\zeta>0$ there exists $\varepsilon^{*}>0$ such that for every $0<\varepsilon<\varepsilon^{*}$, it holds that
+$\int_{\mathbb{R}^{d}}|\mathbf{y}|\mathds{1}_{A_{1/2}^{1}}(\mathbf{y})\mu_{\varepsilon\frac{g(\delta)}{2^{n-1}}}(d\mathbf{y})<m^{*}+\zeta.$
+The last estimate holds for all $n\geq 1$ and for $\delta$ small enough; indeed, since function $g(\delta)$ decreases to 0 monotonically as $\delta\to 0$ it follows that $\varepsilon\frac{g(\delta)}{2^{n-1}}<\varepsilon<\varepsilon^{*}$ for $\delta$ small.
+Consequently, we can deduce that for $\delta$ small, it holds that
+$\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})\leq c_{2}(g(\delta))^{k}\sum_{n=1}^{\infty}2^{-k(n-1)}(m^{*}+\zeta).$
+We can now conclude that (29) holds, which ends the proof by applying Theorem 4.4. ∎
+The next corollary provides a tractable example of the possible measure $\mu$ (and therefore corresponding $ID$ distribution) satisfying the conditions of Lemma 4.5.
 It is a direct generalisation of [GZ24a, Cor.
@@
 Corollary 4.6.
-Let $\varrho_{\varepsilon}$, $\varepsilon>0$ be a collection of $ID(\mathbb{R}^{d})$ measures satisfying (22) and assume that the corresponding Lévy measures $\mu_{\varepsilon}$, $\varepsilon>0$, are of the form
-$\mu_{\varepsilon}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varepsilon\rho(\varepsilon r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}).$
-Assume also that for some $h:\mathbb{S}^{d-1}\mapsto\mathbb{R}$, $h\geq 0$, it holds that
+Consider $\varrho\in ID(\mathbb{R}^{d})$ and assume that its Lévy measure $\mu$, satisfying (27), can be given by
+$\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(\mathbf{x},r)\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
+i.e.
+assume that $\varpi(d\mathbf{x},dr)=\rho(\mathbf{x},r)\sigma(d\mathbf{x})dr$ in (26) for some function $\rho:\mathbb{S}^{d-1}\times(0,\infty)\mapsto[0,\infty)$ and a sigma-finite measure $\sigma$ on $\mathbb{S}^{d-1}$.
+Assume also that for some $h:\mathbb{S}^{d-1}\mapsto\mathbb{R}_{+}$, it holds that
 $\int\limits_{\mathbb{S}^{d-1}}h(\mathbf{x})\sigma(d\mathbf{x})=b<\infty,\quad\lim\limits_{r\downarrow 0}\int\limits_{\mathbb{S}^{d-1}}|r\rho(r,\mathbf{x})-h(\mathbf{x})|\sigma(d\mathbf{x})=0.$
 For simplicity, assume additionally that $b=1$. Then we have
-$\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\tilde{\sigma})\;\text{ as }\varepsilon\downarrow 0,$
-where $\tilde{\sigma}(B)=\int\limits_{B}h(\mathbf{x})\sigma(d\mathbf{x}),B\in\mathscr{B}(\mathbb{S}^{d-1})$.
-Remark 10.
-Let us consider the $ID(\mathbb{R}^{d})$ measure $\varrho$ satisfying (22) with Lévy measure $\tilde{\mu}$ given as
-$\tilde{\mu}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}).$
-For the measure $\tilde{\mu}$ and some $u>0$ we consider the truncation
-$T_{u}\tilde{\mu}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{u}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
-and note that $T_{u}\tilde{\mu}$ is a Lévy measure for any $u>0$.
-Note also that the truncation $T_{u}$ corresponds to considering only small jumps of the underlying Poisson random measure, whereas $\tilde{\mu}-T_{u}\tilde{\mu}$ corresponds to the compound Poisson part.
-Denote then $T_{u}\varrho$ the $ID(\mathbb{R}^{d})$ distribution with Lévy measure $T_{u}\tilde{\mu}$.
-Then $\varrho_{\varepsilon}$ of Corollary 4.5 can be obtained by the transformation $\varrho\mapsto\varepsilon^{-Q}T_{\varepsilon}\varrho$, which accords with truncation and operator rescaling of the original measure.
-Similar constructions have been considered in more detail in [COV09] and [GZ24a].
-The application of this result for simulations has also been discussed there.
+$\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\tilde{\sigma})\;\text{ as }\varepsilon\to 0,$
+where $\tilde{\sigma}(B)=\int\limits_{B}h(\mathbf{x})\sigma(d\mathbf{x})$, $B\in\mathscr{B}(\mathbb{S}^{d-1})$, and $\varrho_{\varepsilon}$ is obtained from $\varrho$ by the transformation described above.
 Note that we can also obtain the following simple corollary from Theorem 4.4, which is an extension of Proposition 3.8iii.
@@
 Proof.
-The proof is analogous to the proof of 3.8iii if we notice that
-$\frac{c_{1}+\cdots+c_{k}}{c}\frac{c_{1}\nu_{1}(\cdot)+\cdots+c_{k}\nu_{k}(\cdot)}{c_{1}+\cdots+c_{k}}\overset{v}{\to}\tilde{\nu}\;\text{as}\;k\to\infty$
-by Lemma 4.3 and use Theorem 4.4. ∎
+The proof is analogous to the proof of Proposition 3.8iii if we notice that
+$\frac{w_{1}\nu_{1}+\cdots+w_{k}\nu_{k}}{w}(D)\begin{cases}\leq\tilde{\nu}(D),&k\in\mathbb{N},\\ \to\tilde{\nu}(D)&\text{as}\;k\to\infty.\end{cases}$
+for all $D\subset\mathbb{S}^{d-1}$, and use Theorem 4.4. ∎
 5 Simulations from the operator Dickman distribution